### v2

# Word problems and embedding-obstructions in cellular automata groups on groups

Abstract
We study groups of reversible cellular automata, or CA groups, on groups. More generally, we consider automorphism groups of subshifts of finite type on groups. It is known that word problems of CA groups on virtually nilpotent groups are in co-NP, and can be co-NP-hard. We show that under the Gap Conjecture of Grigorchuk, their word problems are PSPACE-hard on all other groups. On free and surface groups, we show that they are indeed always in PSPACE. On a group with co-NEXPTIME word problem, CA groups themselves have co-NEXPTIME word problem, and on the lamplighter group (which itself has polynomial-time word problem) we show they can be co-NEXPTIME-hard. We show also two nonembeddability results: the group of cellular automata on a non-cyclic free group does not embed in the group of cellular automata on the integers (this solves a question of Barbieri, Carrasco-Vargas and Rivera-Burgos); and the group of cellular automata in dimension $D$ does not embed in a group of cellular automata in dimension $d$ if $D\geq 3d+2$ (this solves a question of Hochman).
1 Introduction
Reversible cellular automata are the continuous $G$-commuting bijections on $A^{G}$, for a finite alphabet $A$ and a countable group $G$. Equivalently, they are the bijections defined by a local rule (uniform over all $g\in G$), where we only look at cells in a finite neighborhood $gN$ of $g$ to determine the new symbol at $g$, using the local rule. Cellular automata are often seen as a model of parallel computation.
We study the following problem: If we fix a finite set $S$ of cellular automata, how hard is it to determine whether a given word $S^{n}$ gives a nontrivial cellular automaton, when we compose the generators in this order (as functions)? In other words, we study the complexity of the word problem of the group generated by $S$. For finite $S$ we call such groups groups of cellular automata.
In symbolic dynamical terms, $A^{G}$ is a dynamical system called the full shift (on alphabet $A$ and group $G$), and the reversible cellular automata form its (topological) automorphism group. Many of our results are proved for more general sets (subshifts) $X\subset A^{G}$ in place of $A^{G}$.
One may summarize the present paper as follows: if the group $G$ contains subtrees that branch quickly enough (or cones that expand quickly enough), then we can propagate information along these subtrees quickly, and perform fast parallel computation. We obtain several technical statements (and in particular solve two open problems) with this idea, and we now proceed to describe these results.
1.1 Complexity-theoretic results
Appendix A gives the (mostly standard) complexity-theoretic definitions used in the statements.
Our first theorem is barely new, and is stated mainly for context. The hard was shown essentially in [29] (it proves the $\mathbb{Z}$-case) and only a short additional discussion is needed here (see Corollary 8.1). The co-NP algorithm is essentially from [8, 23].
Theorem A.
Let $G$ be an f.g. group with have polynomial growth, and let $X\subset A^{G}$ be any subshift of finite type with language in NP. Then every finitely-generated subgroup of $\mathrm{Aut}(X)$ has word problem in co-NP. Furthermore, for any infinite f.g. group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ has a f.g. subgroup with NP-complete word problem.
For example $X=A^{G}$ satisfies the assumptions on $X$. Growth refers to the size of balls. Recall that polynomial growth corresponds to virtual nilpotency by Gromov’s theorem [17], and that $O(n^{D})$ growth implies $\Theta(n^{d})$ growth for some $d$.
In particular, groups of cellular automata $H\leq\mathrm{Aut}(A^{G})$ always have word problems in co-NP when the group has polynomial growth. Our first main result is about groups with faster growth, namely that if the growth of balls in a group is sufficiently superpolynomial, then the complexity “jumps” from co-NP to PSPACE. It is a standard conjecture in complexity theory that PSPACE is strictly larger than co-NP (equivalently, strictly larger than NP).
Theorem B.
Let $G$ be a f.g. group with at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT where some finite-support point has free shift orbit. Then $\mathrm{Aut}(X)$ has a finitely-generated subgroup with PSPACE-hard word problem.
For example $X=A^{G}$ satisfies the assumptions on $X$. Here, by having at least stretched exponential growth we mean that for some finite generating set (thus for any), the $k$-ball of $G$ has cardinality $\Omega(e^{k^{\beta}})$ for all $k$, for some degree $\beta$.
Under the Gap Conjecture of Grigorchuk [15], the growth rates in the two theorems above are a dichotomy for f.g. groups. Specifically, the version $C^{*}(\beta)$ (for $\beta>0$) of the Gap Conjecture states the following:
Conjecture 1.1.
Every f.g. group that is not of polynomial growth has at least stretched exponential growth with degree $\beta$.
We immediately note that the larger $\beta$ is, the more this conjecture is claiming, and that for large $\beta$ it states too much. Indeed, the conjecture is not true for e.g. $\beta>\log_{32}31=\gamma$, as the Grigorchuk group has growth $O(e^{n^{\gamma}})$ and is not virtually nilpotent [16] (much more is known about the growth of the Grigorchuk group [14]). For our purposes, it would suffice that this conjecture is true for any $\beta>0$.
Grigorchuk has several versions of the Gap Conjecture, and the best-known version is perhaps the one with $\beta=1/2$, but where it is only assumed that $e^{\sqrt{n}}$ is not an upper bound for the growth, rather than that it is a lower bound. For our purposes, it is indeed important to have a stretched exponential lower bound.
Conjecture 1.1 is indeed known to hold in many cases (and for some $\beta>0$): for $\beta=1/6$ it holds for residually solvable and left-orderable groups [15]. It also holds rather trivially (for any $\beta<1$) for many other types of groups, like hyperbolic groups, linear groups and elementary amenable groups. So among these groups, our result shows that the word problem of an automorphism group of a nice enough SFT (e.g. a full shift) is always either in coNP or is PSPACE-hard, and the jump happens exactly after polynomial growth.
In several papers [31, 29, 9, 32], the author (with various coauthors) has studied a specific subgroup of $\mathrm{Aut}(A^{G})$ called $\mathrm{PAut}_{G}(A)$, which is generated by partial shifts (which split the aphabet into a Cartesian product and then shift one of the resulting tracks only), and symbol permutations (permuting the symbols in cells individually). This group is somewhat mysterious, but it provides a reasonably canonical f.g. subgroup, where many behaviors of the typically non-f.g. group $\mathrm{Aut}(A^{G})$ can already be seen.
On a smaller class of groups, we show that this group can be used in the PSPACE-hardness proof.
Theorem C.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet with sufficiently many divisors. Then $\mathrm{PAut}_{G}(A)$ has a PSPACE-hard word problem.
The main idea in all the theorems above is to build permutations of one part of the alphabet which only happen if some circuit evaluates to a true value on the other parts of the alphabet; or in terms of a probabilistic analog, “a particular event happens”. For conditioning on polynomial-support events, this is known as Barrington’s theorem after [3], and it is basically the only ingredient needed in the proof Theorem A in [29].
The main new idea needed is the use of fast information propagation along trees. Specifically, we can also work with events of exponential (or otherwise fast-growing) support, by using the parallelism of cellular automata. Unlike in Barrington’s theorem, it seems we cannot do this while preserving a perfect separation of control and permuted data. The technique we use for this, and in fact in all our constructions, is something we call dyadically accelerated funneling.
For the general result Theorem B (PSPACE-hardness on all groups with fast growth), we have to use some of combinatorics to show that all groups with stretched exponential growth contain large trees in an appropriate sense. This idea originates from a joint work in progress with Ilkka Törmä (though the computational model is different). We also need some symbolic dynamical arguments to construct markers on the nodes of these trees.
In a few cases we show the corresponding upper bound:
Theorem D.
Let $G$ be a free group or a surface group, and let $X\subset A^{G}$ be any subshift of finite type with the finite extension property. Then every finitely generated subgroup of the automorphism group of $X$ has word problem in PSPACE.
For example $X=A^{G}$ satisfies the assumptions on $X$. We give more generally a sufficient geometric condition for such an upper bound in terms of “polynomial splitting schemes” for balls. We leave it as an open question what groups admit such schemes. One interesting question is whether they exist for small cancellation groups or hyperbolic groups with one-dimensional boundary.
Two remarks are in order. First, note that the word problem of $\mathrm{PAut}_{G}(A)$ contains the word problem of $G$, so of course if $G$ itself has PSPACE-hard word problem (or harder) then the theorem is trivial. The word problem is decidable in polynomial time in a large class of groups, and for example for a free group or a surface group it can be solved by Dehn’s algorithm [12, 7].
Second, it is known that automata groups (whose definition we omit) can have PSPACE-complete word problems [36] (and they always have PSPACE word problem [35]). Barrington’s theorem is a key component of the proof in [36], and it would be interesting to know how related the proofs are.
For general groups, the natural upper bound for the complexity of the word problem of the automorphism group of a full shift is not PSPACE, but rather co-NEXPTIME. We show that at least one group (and thus every group containing it) reaches the upper bound.
Theorem E.
On the lamplighter group $\mathbb{Z}_{2}\wr\mathbb{Z}$, for every full shift on at least two letters, the automorphism group has a finitely-generated subgroup with co-NEXPTIME-complete word problem. For any alphabet $A$ with sufficiently many prime factors, we can use the group $\mathrm{PAut}(A)$.
The group $\mathbb{Z}_{2}\wr\mathbb{Z}$ itself has word problem decidable in polynomial time. We believe the analog of Theorem E can be proved similarly on solvable Baumslag-Solitar groups $\mathrm{BS}(1,n)=\mathbb{Z}[\frac{1}{n}]\rtimes\mathbb{Z}$, although we do not show this.
We mention by analog that the domino problem, whether a given SFT is nonempty, is known to be undecidable on $\mathbb{Z}_{2}\wr\mathbb{Z}$, $\mathrm{BS}(1,n)$ [4].
Conjecture 1.2.
Word problems of CA groups can be co-NEXPTIME-complete on a large class of groups.
Some interesting cases to look at would be polycyclic groups (that are not virtually cyclic), hyperbolic groups and $F_{2}\times F_{2}$.
The fact that we show the word problems to be this difficult for “natural” groups may be of interest in itself. There are certainly many general constructions that allow making the word problem of an f.g. group as difficult as one likes while preserving its decidability. A strong result in this direction is [22].
Our results are of a different flavor in that they are simply a chapter in the study of properties of automorphism groups of subshifts, rather than constructed specifically for the purpose of hard word problems.
It seems that in most cases, when the word problem was solved for a group of general interest, it ended up being solvable in polynomial time. In fact, we only know the following cases where the word problem ended up being complete for a class conjecturally not solvable in polynomial time:
  • the higher dimensional Brin-Thompson group $2V$ has co-NP-complete word problem [6];
  • on $\mathbb{Z}$, the groups $\mathrm{PAut}(A)$ have co-NP-complete word problems for suitable alphabets $A$ [29];
  • automata groups can have PSPACE-complete word problems [36].
In the third item, the group was not a specific automata group that was studied before the paper, but rather one constructed for the proof. However, since one such group exists, one can likely find reasonably natural supergroups (generated, for example, by all automata that can be defined with a fixed number of states).
In the second item, the group explicitly used is $\mathrm{PAut}(A)$. One may wonder if this group is really natural: the generators are natural and study of the group emerged from questions unrelated to computability or complexity, but we do not even know, for example, whether it is decidable whether a given $f\in\mathrm{Aut}(A^{\mathbb{Z}})$ belongs to this group. However (unlike in e.g. Theorem C of the present paper), the proof in [29] in fact only uses the subgroup of $\mathrm{PAut}(A,B)$ where the $B$-track cannot affect the $A$-track in symbol permutations, which is a tamer and more natural group in many ways (see e.g. [31]).
A final remark we make here is that in each of the three items above, it is only conjectural (even if widely-believed) that there is no polynomial-time algorithm for the word problem, as PSPACE $\neq$ P is not a proven fact. However, Theorem E unconditionally shows that there is no polynomial-time algorithm for the word problem of $\mathrm{PAut}(A)$ on the lamplighter group, and for alphabets $A$ with sufficiently many prime factors.
1.2 Non-embeddability results
Besides complexity theoretic results, we obtain some non-embeddability results from the same construction.
These are based on an observation that we believe is entirely new (although in retrospect, rather obvious), namely that residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ admits a quantitative variant when working with finitely-generated subgroups, based on the fact that nontrivial automorphism with radius $n$ acts nontrivially on $n$-periodic points. See Lemma 4.1. This idea bears resemblance to the use of simple groups in [34, 13, 32], although we do not know if there is a direct connection.
The following was asked by Hochman in the conference “Current Trends in Dynamical Systems and the Mathematical Legacy of Rufus Bowen” organized in Vancouver in 2017. (A similar question, but allowing an arbitrary SFT in place of $\{0,1\}^{\mathbb{Z}^{d}}$ was asked in [19].)
Question 1.1.
Does $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{D}})$ embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{d}})$ for $d<D$?
We prove the following result:
Theorem F.
Suppose that $D>3d+1$. Then $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any nontrivial alphabets $A,B$.
So, for example, $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{5}})$ does not embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}})$, and $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{8}})$ does not embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{2}})$. There are no previously known nonembeddability results for the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, indeed previously the only embeddability obstructions that have been observed or applied are (qualitative) residual finiteness and complexity of the word problem.
We note that from just $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ not embedding in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$, it would not be obvious that not all finitely-generated subgroups do not embed. For example, it is known that $\mathrm{Aut}(A^{\mathbb{Z}})$ admits a f.g. subgroup that contains an isomorphic copy of every f.g. subgroup of $\mathrm{Aut}(A^{\mathbb{Z}})$, yet it is not known whether there is an f.g. subgroup of $\mathrm{Aut}(A^{\mathbb{Z}})$ that contains an embedded copy of $\mathrm{Aut}(A^{\mathbb{Z}})$ [31].
It should be possible to optimize the proof to get nonembeddings for $D\geq 3d+1$, as our proof “wastes” one dimension, in a sense. It is an interesting question how far this can be pushed, in particular whether all cases $D\geq d+1$ can be covered.
As noted in [2], the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ and $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ are non-isomorphic for $d\neq D$, since their centers are the respective shift groups $\mathbb{Z}^{D}$ and $\mathbb{Z}^{d}$. This is a result of Hochman [19] (generalizing Ryan’s theorem [27], and further generalized by Barbieri, Carrasco-Vargas and Rivera-Burgos in [2]). However, it seems impossible to obtain any results about embeddability using this kind of method.
Hochman asked in personal communication whether Theorem F is also true modulo the centers. Indeed, this is the case, by a minor modification of the proof.
Theorem G.
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
In particular, the groups are not isomorphic modulo their centers.
Next, denote by $F_{2}$ the free group on two generators. The following is a question of Barbieri, Carrasco-Vargas and Rivera-Burgos [2]:
Question 1.2.
Does $\mathrm{Aut}(\{0,1\}^{F_{2}})$ embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}})$?
We solve this question completely.
Theorem H.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any nontrivial alphabet. Then $\mathrm{Aut}(A^{G})$ does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
Note that this is an immediate corollary of Theorem A and Theorem B if PSPACE $\neq$ NP. However, our result is unconditional.
Acknowledgements
We thank Ilkka Törmä for working through the proof of Theorem 4.1 in full detail. Of course, any mistakes are mine. The connection between the Gap Conjecture and a PSPACE frontier originated from the study of another computational model with Ilkka Törmä (apart from the construction of markers). We thank Sebastián Barbieri for sharing Question 1.2, and Scott Schmieding for several inspiring discussions. We thank Laurent Bartholdi for helping streamline the surface group argument (Lemma 8.2), and other useful comments. We thank Michael Hochman for asking about Theorem G.
2 Definitions, conventions and some standard observations
In this section, we explain our definitions and conventions. This does not include complexity-theoretic definitions, which are quite lengthy. We have instead written these in an appendix, which also contains some basic results on complexity theory.
The appendix is aimed mainly at readers with zero knowledge and infinite intelligence, and some readers may find it helpful to also consult a textbook [1]. Readers familiar with complexity theory will not need to read Section A.1 and Section A.2 in the appendix, but may find it helpful to check out Lemma A.3 (an explicit tree formulation of a computation of an alternating Turing machine in terms of two nondeterministic Turing machines). We also give a proof of a version of Barrington’s theorem in Section A.3. A reader familiar with previous work of the author may mostly skip Section A.3, while for others (even readers familiar with Barrington’s theorem) we recommend taking a look, as we use similar notation and arguments in automorphism groups.
2.1 Assorted definitions
[…]
In the context of permutation groups on some set (clear from context), we write the permutation $\pi$ mapping $\pi(a)=b,\pi(b)=c,\pi(c)=a$ and fixing all other elements as $(a\;b\;c)$ (or $(a;\;b;\;c)$ if $a,b,c$ are longer expressions). Such a permutation is called a $3$-rotation.
An alphabet is a finite set of symbols. A word $u$ is an element of the free monoid $A^{*}$ on an alphabet $A$, and its length is the number of $A$-symbols (free generators) it is composed of. We by default 00-index words meaning $u=u_{0}u_{1}\ldots u_{|u|-1}$, but this can be overridden by notation, or if we explicitly say we $1$-index the word i.e. $u=u_{1}u_{2}\ldots u_{|u|}$. We write $A^{\leq n}$ for the words of length at most $n$.
For words $u,v$ we denote the free monoid operation, i.e. concatenation, by $uv$. The empty word (unique word of length 00) is written $\epsilon$, and we write $A^{+}=A^{*}\setminus\{\epsilon\}$. A set of words $W\subset A^{*}$ is mutually unbordered if, whenever $u,v\in W$, $i,j\in\mathbb{Z}$, and $x\in A^{\mathbb{Z}}$ satisfies $x|[i,i+|u|-1]=u$, $x|[j,j+|v|-1]=v$, we have $[i,i+|u|-1]\cap[j,j+|v|-1]=\emptyset$, unless $i=j\wedge u=v$. A word $w$ is unbordered if $\{w\}$ is mutually unbordered.
We use the usual notations $O(f(n)),\Theta(f(n)),\Omega(f(n))$: $f^{\prime}(n)=O(f(n))$ if $f^{\prime}(n)\leq Cf(n)$ for large enough $n$ and some $C>0$; $f^{\prime}(n)=\Omega(f(n))$ if $f^{\prime}(n)\geq Cf(n)$ for large enough $n$ and some $C>0$; $f^{\prime}(n)=\Theta(f(n))$ if both are true.
Vertices of a graph of any kind are written $V(G)$ and edges as $E(G)$. We consider undirected graphs whose edges are sets $\{u,v\}$ of cardinality $2$ with $u,v\in V(G)$, as well as directed graphs (defined later).
In a metric space $X$, we say $Y\subset X$ is $r$-dense if every $r$-ball intersects $Y$. We say $Y$ is $r$-sepated if $d(y,y^{\prime})<r$ for $y,y^{\prime}\in Y$ implies $y=y^{\prime}$. The following is immediate in finite metric spaces (in general, axiom of choice is needed):
Lemma 2.1 (Packing Lemma).
In any metric space $X$, for any $r$ there exists an $r$-dense $r$-separated subset $Y\subset X$.
Proof.
If $Y$ is maximal $r$-separated, then we cannot add any new points, so for every $x\in X$, $\{x\}\cup Y$ is not $r$-separated, i.e. $d(x,y)<r$ for some $y\in Y$. We conclude that any such maximal $Y$ is $r$-dense. ∎
2.2 Groups
Our groups $G$ always have a fixed generating set $S$ which is symmetric $S=\{s^{-1}\;|\;s\in S\}$ and $1_{G}\in G$, and $G$ is considered as a metric space with the corresponding word metric
$d(g,h)=\min\{n\;|\;h=gs_{1}s_{2}\ldots s_{n}\mbox{ for some }s_{i}\in S\},$
and ($k$-)balls are $B_{k}=B_{k}^{G}=\{g\in G\;|\;d(1_{G},G)\leq n\}$ or their translates $gB_{k}$. A geodesic in a group is a sequence of group elements $g_{1},\ldots,g_{n}$ such that $d(g_{i},g_{j})=|i-j|$ for $i,j$. It is obvious that geodesics of arbitrary length exist in f.g. groups.
The commutator of group elements $a,b$ is $[a,b]=a^{-1}b{-1}ab$; $[G,G]$ is the group generated by commutators $[a,b]$ where $a,b\in G$. Also define longer commutators by $[a_{1},a_{2},\ldots,a_{k}]=[a_{1},[a_{2},\ldots,a_{k}]]$. A group is abelian if any two elements $g,h$ commute. A group $G$ is nilpotent if the lower central series $G,G_{1}=[G,G],G_{2}=[G,G_{1}],G_{3}=[G,G_{2}],\ldots$ reaches the trivial group in finitely many steps. A group $G$ is solvable if the derived series $G,G^{1}=[G,G],G^{2}=[G^{1},G^{1}],\ldots$ reaches the trivial group in finitely many steps. The center of a group $G$ is the set of elements $g\in G$ that commute with all other elements.
Two groups $G,H$ are isomorphic if there is a bijection between them that preserves the multiplication relation, and we write this as $G\cong H$. The groups $G,H$ are commensurable if there exist subgroups $L\leq G,L^{\prime}\leq H$ of finite index such that $L\cong L^{\prime}$. The index $[G:H]$ ofa subgroup $H\leq G$ is the cardinality of $G/H=\{gH\;|\;g\in G\}$.
A group $G$ is virtually P (where P is any group, property of groups or family of groups) if $G$ has a subgroup of finite index which is P. It is residually P if for all $g\in G$ there is a quotient in $P$ where $g$ maps nontrivially. A group $G$ is amenable if for all $\epsilon>0$ and $S\subset G$ finite, there exists finite $F\subset G$ such that $\frac{FS}{F}<1+\epsilon$. A group $G$ is left-orderable if there exists a total order $<$ on $G$ such that $a<b\implies ga<gb$ for all $a,b,g\in G$.
The symmetric group (of all permutations, under composition) on a finite set $A$ is $\mathrm{Sym}(A)$, and $\mathrm{Alt}(A)$ is the simple index-$2$ subgroup of even permutations. For integer $n$, we write $\mathrm{Sym}(n),\mathrm{Alt}(n)$ for $\mathrm{Sym}([1,n]),\mathrm{Alt}([1,n])$ respectively.
2.3 Word problems
Now we describe our word problem conventions. We will work with word problems of groups that are not finitely-generated, but are mainly interested in their finitely-generated subgroups. Thus, we use a non-standard convention for word problems. We define $\mathrm{WPs}(G)$ to be the set of word problems of its finitely-generated subgroups. This is a family of languages, and we next define our conventions for complexity theoretic notions of language families.
If $C$ is a complexity class of languages defined directly by resource bounds (such as the classes P, co-NP, PSPACE or co-NEXPSPACE), then for a family of languages $\mathcal{L}$, we say that $\mathcal{L}$ is in $C$ if every language in $\mathcal{L}$ is in $C$. We say $\mathcal{L}$ is $C$-hard if some language in $\mathcal{L}$ is $C$-hard (for polynomial time many-one reductions). We say $\mathcal{L}$ is $C$-complete if $\mathcal{L}$ is in $C$ and is $C$-hard.
Note that $D$-hard (and $D$-complete) can themselves be called complexity classes, and we do not give a definition of “resource bounded” that excludes such classes. Thus saying a language class $\mathcal{L}$ is $D$-hard can seem ambiguous. This will not cause problems; simply do not think of $D$-hard and $D$-complete as complexity classes, but instead derived notions that are defined differently for languages and language families.
2.4 Symbolic dynamics
In the context of a group $G$ and alphabet $A$, a pattern is $p\in A^{D}$. If $q\in A^{E}$ for some $E\subset G$, $q$ contains a translate of $p$ if there exists $g\in G$ such that $\forall h\in D:q_{gh}=p_{h}$.
The set $A^{G}$ for finite alphabet $A$ and countable group $G$ is given its product topology (making it a Cantor space), and it has a natural continuous action of $G$ by $gx_{h}=x_{g^{-1}h}$ for $g,h\in G,x\in A^{G}$. The full shift is $A^{G}$ as a dynamical system under the $G$-action. A subshift is a $G$-invariant closed subset of $A^{G}$ (i.e. a subsystem of a full shift). A subshift of finite type or SFT is $X\subset A^{G}$, such that for some clopen set $C\subset A^{G}$, we have $X=\{x\in A^{G}\;|\;\forall g\in G:gx\notin C\}$ (such a set is indeed a subshift).
A subshift has the finite extension property, or is FEP if there exists a finite set $D\subset G$, a set of patterns $P\subset A^{D}$ and a look-ahead $r\in\mathbb{N}$ such that for all $E\subset G$ and $x\in A^{E}$, we have $x\in X|E$ (i.e. $x$ is globally valid) if and only if there exists $y\in A^{EB_{r}}$ such that $y$ does not contain any translate of a pattern from $P$ (i.e. $y$ is locally valid). An FEP subshift is always an SFT. Generalizing FEP, we say a subshift has the polynomial extension property for balls or PEPB if, with respect to some finite set of forbidden patterns $P$, and for some degree $t$, a pattern $p\in A^{B_{n}}$ is globally valid if and only if some extension $q\in A^{B_{n^{t}}}$ is locally valid.
The elements $x\in A^{G}$ are called configurations or points. The support of a configuration is $\{g\in G\;|\;x_{g}\neq 0\}$, where $0\in A$ is some agreed-upon zero-symbol (usually precisely 00).
An endomorphism of a subshift $X$ is a $G$-equivariant continuous $\phi:X\to X$. If there is a two-sided inverse (which is also an endomorphism), $\phi$ is an automorphism. In the case $X=A^{G}$, endomorphisms are called cellular automata and automorphisms reversible cellular automata. An endomorphism always has a radius $r\in\mathbb{N}$ and a local rule $f_{\mathrm{loc}}:A^{B_{r}}\to A$ such that $f(x)_{g}=f_{\mathrm{loc}}(x|gB_{r})$, where we see $x|gB_{r}$ as an element of $A^{B_{r}}$ in the obvious way.
Two subshifts $X\subset A^{G},Y\subset B^{G}$ are (topologically) conjugate if there is a homeomorphism $\phi:X\to Y$ that commutes with the $G$-actions; $\phi$ is called a conjugacy and it conjugates $X$ with $Y$. Generalizing endomorphisms and conjugacies, a continuous shift-commuting map $\phi:X\to Y$ is called a block map. When such a map is injective, it conjugates $X$ with its image. The SFT property is known to be preserved under conjugacy. Also block maps (and thus conjugacies) have local rules, defined in an analogous way as for cellular automata.
2.5 $\mathrm{PAut}$
We define a reasonably natural (or at least canonical) f.g. subgroup of the automorphism group that we use in some of the results.
Definition 2.1.
Let $G$ be a group, and let $A_{1},...,A_{k}$ be finite sets and $A=\prod_{i}A_{i}$. Then $\mathrm{PAut}_{G}(A_{1},...,A_{k})$ is the group generated by
  • partial shifts $\sigma_{g,i}$ for $g\in G,i\in[1,k]$, which on input $x=(x_{1},\ldots,x_{k})\in A^{G}$ map $\sigma_{g,i}(x)=(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{k})$ where $y_{h}=x_{hg}$ for $h\in G$; and
  • symbol permutations $\eta_{\pi}$ for $\pi\in\mathrm{Sym}(A)$, which on input $x\in A^{G}$ are defined by $f_{\pi}(x)_{h}=\pi(x_{h})$ for $h\in G$.
For any finite alphabet $A$, define $\mathrm{PAut}(A)$ to be $\mathrm{PAut}(A_{1},...,A_{k})$ where $A\cong A_{1}\times\cdots\times A_{k}$ through any bijection, where $|A_{i}|$ is prime.
When $G$ is clear from context, we write just $\mathrm{PAut}(A)=\mathrm{PAut}_{G}(A)$, or even just $\mathrm{PAut}$ if also the alphabet is clear from context.
Lemma 2.2.
$\mathrm{PAut}(A)$ is well-defined.
Proof.
This is entirely analogous to [31, Lemma 2.1]. The point is the change between the groups for different choices is in the partial shifts. Since the prime decomposition is unique, these are conjugated by a symbol permutation, which is in the group in any case. ∎
3 Groups with free submonoids
In this section, we prove PSPACE-hardness of the word problem for full shifts with suitable alphabets on groups that contain free submonoids on at least two generators.
Theorem 3.1.
Let $G$ be a group containing a free monoid on two generators. Then for some $A$, the group $\mathrm{PAut}(A)$ has PSPACE-hard word problem.
Proof.
If $G$ contains a free monoid on two generators, it contains a free monoid on three generators, say powers of $a,b,c\in G$ satisfy no nontrivial positive relations.
We will use alphabet $A=\mathbb{Z}_{7}\times B\times\{0,1,2\}$ for our f.g. group, where $B=Q\times S$. Here, $Q$ with $|Q|=t$, and $S$ with $|S|=s$ are respectively the common sets of states and tape symbols of the nondeterministic Turing machines $T_{1},T_{2}$ from Lemma A.3.
Our reduction takes a word $u_{\mathrm{init}}$ of length $n=2^{k}$ (note that one should think of this as a polynomial size input whose size just happens to be of nice form, rather than an exponential length word), and constructs an automorphism that performs a permutation $\pi$ on the $\mathbb{Z}_{7}$-component of $A$ whenever it sees a tree of the form described in Lemma A.3 on the configuration, encoded in words read on the $B$-components, and with $q_{\mathrm{init}}u_{\mathrm{init}}\in QS^{n}$ as the initial configuration on the tree (where $q_{\mathrm{init}}$ is the common initial state of $T_{1},T_{2}$).
As in the lemma, the computation steps in the tree are given by the Turing machines $T_{0},T_{1}$ with common initial state $q_{\mathrm{init}}$, states $Q$ and tape alphabet $S$. Note that we now have the state serially next to tape symbols rather than use a Cartesian product, just for notational simplicity. The tree still has depth $n$, but the Turing machine configurations on it have length $n+1$.
This will indeed many-one reduce the language $L$ from Lemma A.3 to the word problem of the automorphism group of $A^{G}$, since the automorphism corresponding to the word $u_{\mathrm{init}}$ will be nontrivial if and only if the permutation is ever performed, and it is performed if and only if a tree proving $u_{\mathrm{init}}\in L$ exists.
The main difficulty here is that the tree has exponential size in $|u_{\mathrm{init}}|$, so we cannot use Barrington’s theorem to directly verify the shape of the tree (which is essentially all we had to do in the $\mathbb{Z}$-case in [29]). The computation will be “initialized” with Barrington’s theorem, and in addition we will use a tree propagation argument to efficiently send a permutation down the computation tree in order to check that the computation is correct everywhere.
We want to use the following generators, which are obviously in $\mathrm{Aut}(A^{G})$:
  • partial right shifts of the first track, i.e. writing $x\in A^{G}$ as $(x_{1},x_{2})\in\mathbb{Z}_{7}^{G}\times(\mathbb{Z}_{m}\times\{0,1\})^{G}$, we take the map $\hat{g}(x_{1},x_{2})_{c}=((x_{1})_{cg},(x_{2})_{c})$ for all $g\in G$.
  • (even) symbol permutations affecting only $\mathbb{Z}_{7}$: for each $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ and each $s\in\mathbb{Z}_{m}\times\{0,1\}$, and writing $x\in A^{G}$ as $(x_{1},x_{2})$ as above, we take the permutation $\alpha_{\pi|s}$, defined by
    $\alpha_{\pi|s}(x_{1},x_{2})_{c}=\begin{cases}(\pi(x_{1})_{c},(x_{2})_{c})&\mbox{ if }(x_{2})_{c}=s\\ ((x_{1})_{c},(x_{2})_{c})&\mbox{ otherwise.}\end{cases}$
  • for each permutation $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$, $s\in\mathbb{Z}_{7}$ not in the support of $\pi$ and $g\in G\setminus\{1\}$, the map $\phi=\phi_{\pi,s,g}$ defined by
    $\phi(x_{1},x_{2})_{c}=\begin{cases}(\pi((x_{1})_{c}),(x_{2})_{c})&\mbox{ if }(x_{1})_{cg}=s\\ ((x_{1})_{c},(x_{2})_{c})&\mbox{ otherwise.}\end{cases}$
Note that the generators only affect the first track (though the way it is changed can depend on the latter tracks). Call this group $\mathcal{G}$.
We show that $\mathcal{G}$ is a subgroup of $\mathrm{PAut}=\mathrm{PAut}(\mathbb{Z}_{7},B,\{0,1\})$, as required in the statement. This is only nontrivial for $\phi_{\pi,s,g}$. To show it is in $\mathrm{PAut}$, we will use the $\{0,1,2\}$-component of the state to “move” a permutation from one cell to another. The $\{0,1,2\}$-component acts here as a borrowed bit in the terminology [37] (or perhaps borrowed trit would be more accurate); this means that while it is used in the construction, its initial value is irrelevant and this value is returned at the end.
We now give the detailed proof of $\phi_{\pi,s,g}\in\mathrm{PAut}$. Write $\pi=[\pi_{0},\pi_{1}]$ where neither $\pi_{i}$ has $s$ in its support. We define four automorphisms by describing their effect on the symbol $x_{c}$ of the input $x\in A^{G}$, for $c\in G$. Define an automorphism $\theta_{b}$ that, if the $\{0,1,2\}$-component of the alphabet at $x_{c}$ is $b\in\{0,1\}$, permutes the first component of $x_{c}$ by $\pi_{i}$. Clearly this is in $\mathrm{PAut}$. Then define an automorphism $\theta$ that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) if the $\mathbb{Z}_{7}$-component of $x_{cg}$ is $s$. This is conjugate to a symbol permutation by a partial shift, so again in $\mathrm{PAut}$. Finally, let $\theta^{\prime}$ be the automorphism that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) in any case. Note that
We claim that
[…]
From now on, let us write configurations $x\in A^{G}$ as $x=(x_{1},x_{2},x_{3})\in\mathbb{Z}_{7}^{G}\times\mathbb{Z}_{m}^{G}\times\{0,1\}^{G}$. The TM configuration at $g$ is $i\mapsto x_{2}|gc^{i}$ for $i\in\{0,\ldots,n\}$. Note that is a word in $(S\times Q)^{n+1}$, and might not actually be a TM configuration in the usual sense (as it might contain multiple states, or none). We say a cell $g$ is TM-good if the TM configuration at $g$ has exactly one symbol in $Q$.
The index at $g$ is $i\mapsto x_{3}|gc^{i}$ for $i\in\{0,\ldots,n-1\}$. It is a ternary word of length $n$ over alphabet $\{0,1,2\}$, which we interpret as a index to the tree witness from Lemma A.3. The symbol $2$ is used for padding the words. If the index at $g$ is $uv$ where $u\in\{0,1\}^{*}$ and $v\in 2^{*}$, then the counter value at $g$ is $|u|$. Note that the counter value may not exist for all $g$.
(As an aside, in the present proof, only the counter value at $g$ really matters in the following, so we could equally well use 00 in place of 00 and $1$, and $1$ in place of $2$, but above it was useful to have three states in this component, and in later variants of this argument we will have more uses for the precise indices.)
We say $g$ is good (with respect to a configuration $x\in A^{G}$) if
  • $g,ga,gb$ are TM-good,
  • the TM-configurations $g$ and $ga$ are in the transition relation of $T_{1}$, and the ones at $g$ and $gb$ are in the transition relation of $T_{2}$,
  • if the index at $g$ is in $u2^{*}$ for $u\in\{0,1\}^{*}$, the indices at $ga,gb$ are in $u02^{*},u12^{*}$ respectively,
  • if the counter value at $g$ is 00, then the TM configuration is precisely $q_{0}u_{\mathrm{init}}$.
[…]
Let $\gamma_{\pi,\ell}$ be the automorphism that permutes $\mathbb{Z}_{7}$ at $g$ by $\pi$ if and only if the counter value is exactly $\ell$. Again, one easily checks that there is a logarithmic depth formula for comparing a number to a constant, and uses Theorem A.2.
Next, we perform what we call dyadically accelerated funneling which we shorten as DAF. In such a proof, a permutation will intuitively travel down a “funnel” of depth $n$. In the present proof, the funnel is a full binary tree, but we also use other kinds of trees (in Theorem 5.1) and discrete analogs of Euclidean cones (Theorem 4.1) in other proofs. [^1] We introduce a term (DAF) for this (informal!) concept, so we can refer back to this idea in other proofs, and since we consider this the main new idea of the paper as far as constructions go. To explain the terminology a little, our mental picture of “funneling” is that “verification checks” travel down a funnel in the form of controlled permutations in intermediate automorphisms in the construction, and finally a permutation is extracted at the bottom. Then the point is that by an induction on dyadic intervals, we can perform such a funneling process in logarithmically many steps.
For $i\in[0,k]$ and $j\in[0,2^{k-i}-1]$, consider now an interval $I=[j2^{i},(j+1)2^{i}]\subset[0,n]$, let $s\in\mathbb{Z}_{7}$ and $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ is a $3$-rotation. Note that $|I|\geq 2$ always. We define an automorphism $\psi_{j,i,s,\pi}$ as follows: The latter two tracks never change their values. On the first track, in position $c\in G$, the value changes by an application of $\pi$ if and only if the following hold:
  • the counter value at $c$ (exists and) is equal to $j2^{i}$;
  • for all $\ell\in[0,2^{i}-1]$ and any $h\in\{a,b\}^{\ell}$, the node $ch$ is good;
  • for all $\ell\in[1,2^{i}-1]$ and any $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is 00;
  • for all $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is $s$.
Note that the counter value at $ch$ is $j2^{i}+\ell$ for all $\ell\in[0,2^{i}]$ and any $h\in\{a,b\}^{\ell}$ by the first two items.
The DAF argument is specifically the proof (method) showing that the automorphisms $\psi_{0,k,s,\pi}$ have polynomial word norm in $n$ in the group $\mathcal{G}$. Note that this is not immediate from Barrington’s theorem, from which for checking conditions on a tree with size exponential in $n$, we can only conclude a word norm polynomial in $2^{n}$.
Suppose first that $s$ is not in the support of $\pi$. Then for $i=0$ and $j\in[0,2^{k-i}-1]$ we can build the above automorphisms from the automorphisms $\beta_{\pi}$, $\phi_{\pi,s,g}$ and $\gamma_{\pi,\ell}$ described above. Specifically, let $\pi=[\pi_{0},\pi_{1},\pi_{2},\pi_{3}]$ where the $\pi_{i}$ are $3$-rotations not having $s$ in their support. We claim that
$\psi_{j,0,s,\pi}=[\beta_{\pi_{0}},\phi_{\pi_{1},s,a},\phi_{\pi_{2},s,b},\gamma_{\pi_{3},j}].$
The norm is clearly polynomial in $n$.
As for correctness of the formula, the logic is similar to Lemma A.9. Namely, the four automorphisms that appear in the commutator only change the $\mathbb{Z}_{7}$-values, and do not change the positions where the $s$ appears in the $\mathbb{Z}_{7}$-component never change. Thus in each cell, the permutation $\pi_{i}$ (or its inverse) either applies each time the corresponding automorphism is applied, or never applies.
We conclude that the only way their effect does not cancel in a particular node is that all of them perform the permutation in that node $h$. That $\beta_{\pi_{0}}$ is performed means $h$ is good. That $\phi_{\pi_{1},s,a}$ and $\phi_{\pi_{1},s,b}$ perform the permutation means $ha$ and $hb$ both contain $s$ in the $\mathbb{Z}_{7}$-component, and that $\gamma_{\pi_{3},j}$ is performed means that the counter value is $j$ at $h$. This is precisely the set of conditions under which $\pi=[\pi_{0},\pi_{1},\pi_{2},\pi_{3}]$ is supposed to be performed, proving the claimed equality for $i=0$ and $s$ not in the support of $\pi$.
[…]
still has norm polynomial in $n$.
[…]
Let $N_{0}$ be an upper bound on the norms of the automorphisms $\psi_{j,0,s,\pi}$, which we have now concluded to be polynomial in $n$, say $O(n^{t})$.
The crux of DAF is to perform induction to get $\psi_{j,i,s,\pi}$ for larger $i$, with only exponential blow-up as $i$ is increased. We have
$[j2^{i},(j+1)2^{i}]=[2j2^{i-1},(2j+1)2^{i-1}]\cup[(2j+1)2^{i-1},2(j+1)2^{i-1}]$
and the two intervals have intersection $\{(2j+1)2^{i-1}\}$. We claim that
[…]
where $\pi=[\pi_{1},\pi_{2}]$, and $\pi_{3}$ is any permutation taking 00 to $1$.
[…]
Suppose then that $\psi_{j,i,s,\pi}$ does not apply the permutation. There are several reasons why this might not happen:
  • we may have that for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$; or
  • for some $\ell\in[1,2^{i}-1]$ and some $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is not 00; or
  • for some $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is not $s$.
We look at these three cases one by one. In the first case, for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$. Since none of the three automorphisms in the right-hand-side formula change the counter values, this will stay true at $c$ at all the applications. Thus, depending on whether $\ell\in[0,2^{i-1}]$ or $\ell\in[2^{i-1},2^{i}]$, either the maps $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ will not apply, or the conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ will not apply. In the former case, nothing happens, and in the latter case, $\psi_{2j+1,i-1,s,\pi_{3}}$ does not flip one of the values in $c\{a,b\}^{2^{i-1}}$ meaning one of $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ ends up not applying a permutation at $c$ (because for them both to apply, the entire frontier must flip its value), which leads to the permutation cancelling at $c$.
Suppose then that we are not in the first case, so the counter value at $ch$ is equal to $j2^{i}+\ell$ whenever $\ell\in[0,2^{i}]$ and $h\in\{a,b\}^{\ell}$. Suppose that we are, however, in the third case, so for some $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is not $s$. Then, $\psi_{2j+1,i-1,s,\pi_{3}}$ never applies, and we are in a similar case as above.
Suppose then that we are not in the first or third case, but that we are in the second case. So for some $\ell\in[1,2^{i}-1]$ and some $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is not 00. If this happens for some $\ell\neq 2^{i-1}$, we are in a similar situation as in the previous case: either $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ will not apply, or the conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ will not apply, and everything cancels at $c$.
If this only happens with $\ell=2^{i-1}$, everything is as in the case where $\psi_{j,i,s,\pi}$ would normally apply, except that at exactly depth $2^{i-1}$, some values are not necessarily 00. The conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ now always applies. Nevertheless, again one of $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ ends up not applying a permutation at $c$.
This case analysis concludes the proof of Equation 2.
The word norm of such an element at level $i$ is now $N_{i}\leq 8N_{i-1}$, so $N_{k}\leq 8^{\log_{2}2^{k}}=8^{k+1}N_{0}=O(n^{3}n^{t})$.
Thus, $\psi_{0,k,s,\pi}$ can be presented as a word over the generators of $\mathrm{PAut}$ with word norm polynomial in $n$ (which in turn, as we recall, is the length of the initial configuration $u_{\mathrm{init}}$). As discussed in the beginning of the proof, this element is nontrivial if and only if there is a way to fill a tree with a computation of $T_{1}$ and $T_{2}$ starting from $q_{0}u_{\mathrm{init}}$, concluding the reduction from $L$ to the word problem of $G$ (thus $\mathrm{PAut}$). ∎
4 Unconditional nonembedding results
The following lemma is the basis of our nonembedding results. It is a simple way to express that the size of quotients needed in the definition of residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ has an upper bound in terms of the radii of automorphisms.
Lemma 4.1.
Let $A$ be a finite alphabet, let $\mathcal{G}=\langle S\rangle\leq\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, where $S$ is a finite generating set defining the word norm of $\mathcal{G}$. Then there exists $C\in\mathbb{N}$ such that whenever $H\leq\mathcal{G}$ is a simple group containing a nontrivial element of norm at most $n$, we have $H\leq\mathrm{Sym}(C^{n^{d}})$.
In particular, $H\leq\mathrm{Sym}(C^{n^{d}})$ implies that $H$ has cardinality at most $|C|^{n^{d}}!$.
Proof.
Let $r$ be the maximal radius among generators in $S$. Suppose $h\in H$ has norm $n$. Then the radius of $h$ is at most $rn$. Since $h$ is nontrivial, the local rule of $h$ acts nontrivially on some pattern $p\in A^{B_{rn}}$. Thus, $h$ acts nontrivially on the set of points with period $(2r+1)n$ in each direction (since $(2r+1)n\geq 2rn+1$ and $B_{rn}\subset[-rn,rn]^{d}$).
Thus, $h\in H$ has nontrivial image in the natural quotient $H\to\mathrm{Sym}(A^{((2r+1)n)^{d}})$ arising from the action on periodic points of such period. Since $H$ is simple, and the kernel of this map is not $H$, the kernel must be trivial, so $H$ embeds in $\mathrm{Sym}(|A|^{((2r+1)n)^{d}})=\mathrm{Sym}(C^{n^{d}})$ where $C=|A|^{(2r+1)^{d}}$. ∎
Note that if $H$ has at least three elements, then $H$ must embed in the simple group $\mathrm{Alt}(C^{n^{d}})$.
Theorem 4.1.
Suppose that $D>3d+1$, and let $A,B$ be nontrivial finite alphabets. Then $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$.
Proof.
First, consider the case $A=\mathbb{Z}_{7}\times\{0,1\}^{D-1}$. In this case, we show that $\mathrm{PAut}(A)$ can be used in the proof.
For the following discussion, fix $n=2^{k}\in\mathbb{N}$ and a configuration $x\in A^{\mathbb{Z}^{D}}$. The counters of $\vec{v}$, written $C(\vec{v})$ are the $D-1$ many numbers whose binary representations of length $k+1$ are written on the $\{0,1\}^{D-1}$-component (each cell contains one bit of each number), in the consecutive positions $(\vec{v},\vec{v}+e_{D},\ldots,\vec{v}+ne_{D})$. The components in this vector must all have a value in $[0,n]=[0,2^{k}]$, or otherwise the counters are not defined. A coordinate $\vec{v}$ is good if its counters $\vec{u}$ exist, and for each natural generator $e_{i}$ of $\mathbb{Z}^{D}$ except $e_{D}$, the counters of $\vec{v}+e_{i}$ exist and are $\vec{u}+e_{i}$.
We define analogs of automorphisms used in the proof of Theorem 3.1. Let $\pi$ be an even permutation of $\mathbb{Z}_{7}$, and define $\beta_{\pi}$ as the automorphism that performs $\pi$ on $\mathbb{Z}_{7}$ at $\vec{v}$ if the coordinate $\vec{v}$ is good. From Lemma A.15 (though the input is formatted slightly differently) we see that $\beta_{\pi}$ has word length polynomial in $n$ in the word norm of $\mathrm{PAut}(A)$.
For $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ and $s$ not in the support of $\pi$, let $\phi_{\pi,s,i}$ be the automorphism that performs $\pi$ on $\mathbb{Z}_{7}$ at $\vec{v}$ if the $\mathbb{Z}_{7}$-value at $\vec{v}+e_{i}$ is $s$. Analogously to Equation 1 in the proof of Theorem 3.1, we see $\phi_{\pi,s,i}\in\mathrm{PAut}(A)$ with constant word norm (instead of the borrowed trit in the $\{0,1,2\}$-component, one can use one of the bits in the $\{0,1\}^{D-1}$-component as a borrowed bit that returns to its original value after the operation). Let $\gamma_{\pi,\ell}$ be the automorphism that permutes $\mathbb{Z}_{7}$ at $\vec{v}$ by $\pi$ if and only if the sum of the counters at $\vec{v}$ is exactly $\ell$. This has polynomial word norm by Lemma A.16 (again, understanding that inputs are formatted slightly differently).
Define $S_{i}=\{\vec{v}\;|\;\vec{v}\geq 0\wedge\vec{v}_{D}=0\wedge\sum\vec{v}=i\}$, a discrete analog of the standard simplex, and define $T_{i}=\bigcup_{j\leq i}S_{j}$. Note that $|T_{i}|=\sum_{j=0}^{i}\geq C^{\prime}i^{D-1}$ for some global constant $C^{\prime}$. Note also that $\gamma_{\pi,\ell}$ acts nontrivially precisely when the counter value is $\vec{v}$ such that $\vec{v}^{\prime}\in S_{\ell}$, where $\vec{v}^{\prime}_{i}=\vec{v}_{i}$ for $1\leq i<D$ and $\vec{v}^{\prime}_{D}=0$. In what follows, we often identify vectors of dimension $D-1$ with vectors of dimension $D$ with 00 in the last coordinate.
We now perform dyadically accelerated funneling along the sets $T_{i}$, analogously to the proof of Theorem 3.1. For $i\in[0,k]$ and $j\in[0,2^{k-i}-1]$, let $I=[j2^{i},(j+1)2^{i}]\subset[0,n]$, let $s\in\mathbb{Z}_{7}$ and $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ a $3$-rotation. Note that $|I|\geq 2$ always.
We define an automorphism $\psi_{j,i,s,\pi}$ as follows for $\pi$ a $3$-rotation: The $\{0,1\}^{D-1}$-components of cells never change their values. On the first track, at $\vec{v}$, the value changes by an application of $\pi$ if and only if the following hold:
  • the counters $C(\vec{v})$ exists and $\sum C(\vec{v})=j2^{i}$ i.e. $C(\vec{v})\in S_{j2^{i}}$;
  • for all $\ell\in[0,2^{i}-1]$ and any $\vec{u}\in S_{\ell}$, the node $\vec{v}+\vec{u}$ is good;
  • for all $\ell\in[1,2^{i}-1]$ and any $\vec{u}\in S_{\ell}$, the $\mathbb{Z}_{7}$-value at $\vec{v}+\vec{u}$ is 00;
  • for all $\vec{u}\in S_{2^{i}}$, the $\mathbb{Z}_{7}$-value at $\vec{v}+\vec{u}$ is $s$.
Note that by the first two items, the counter value $C(\vec{v}+\vec{u})=j2^{i}+\ell$ for all $\ell\in[0,2^{i}]$ and any $\vec{u}\in S_{\ell}$.
Suppose first that $s$ is not in the support of $\pi$. Then for $i=0$ and $j\in[0,2^{k-i}-1]$ we can build the above automorphisms from the automorphisms $\beta_{\pi}$ and $\phi_{\pi_{1},s,1}$ described above: by letting $\pi=[\pi_{0},\pi_{1},\ldots,\pi_{D+1}]$ we have
$\psi_{j,0,s,\pi}=[\beta_{\pi_{0}},\phi_{\pi_{1},s,1},\ldots,\phi_{\pi_{D},s,D-1},\gamma_{\pi_{D+1},j}].$
The norm is polynomial in $k$ (though there is a multiplicative constant exponential in $D$).
The idea for the proof of correctness is that we can interpret the maps
$\beta_{\pi_{0}},\phi_{\pi_{1},s,1},\ldots,\phi_{\pi_{D},s,D-1},\gamma_{\pi_{D+1},j}$
as controlled permutations at a given node $\vec{v}$ checking the various requirements for $\psi_{j,0,s,\pi}$ applying, and as in Lemma A.9 the effect of the iterated commutator is to perform the iterated commutator of the $\pi_{i}$ if all these conditions hold, and otherwise do nothing. The detailed proof is exactly analogous to the corresponding claim in Theorem 3.1, but using $T_{j}$ in place of the tree $\{a,b\}^{\leq 2^{j}}$ (unlike in that proof, the shifts of $T_{j}$ appearing in the proof have nontrivial intersections, but this does not affect the logic in any way).
[…]
still has norm polynomial in $k$. Again, the detailed proof is as in Theorem 3.1. Let $N_{0}$ be an upper bound on norms of the maps $\psi_{j,0,s,\pi}$, which we have concluded is polynomial in $k$, i.e. it is $O(\log(n)^{t})$ for some $t$.
We perform induction on $i=0,\ldots,k$ to get the maps $\psi_{j,i,s,\pi}$. We have
$[j2^{i},(j+1)2^{i}]=[2j2^{i-1},(2j+1)2^{i-1}]\cup[(2j+1)2^{i-1},2(j+1)2^{i-1}]$
and the two intervals have intersection $\{(2j+1)2^{i-1}\}$. It is easy to check (see again the proof of the corresponding claim in the proof of Theorem 3.1) that
[…]
where $\pi=[\pi_{1},\pi_{2}]$, and $\pi_{3}$ is any permutation taking 00 to $1$.
Writing $N_{i}$ for the maximal word norm of such an element $\psi_{j,i,s,\pi}$, the formula shows $N_{i}\leq 8N_{i-1}$, so $N_{k}\leq 8^{k}N_{0}=O(n^{3}\log(n)^{t})$ for some $t$.
The automorphism $\psi_{0,n,\pi,0}$ can now fully described as follows: Say $\vec{v}$ has a marker (and $\vec{v}$ is the root of the marker) if for all $\vec{u}\in T_{n}$, the counter values in $\vec{v}+\vec{u}$ are $\vec{u}$. Note that this provides a full description of the contents of $\vec{v}+\vec{u}+\vec{w}$ for $\vec{w}\in[0,k]e_{D},\vec{u}\in T_{n}$). Enumerate the $M=\Omega(n^{D-1})$ vectors in $T_{n}$ as $\vec{t}_{1},\ldots,\vec{t}_{M}$, starting with $\vec{t}_{1}=\vec{0}$. Then $\psi_{0,n,\pi,0}$ is the automorphism that, whenever $\vec{v}$ has a marker, performs the $3$-rotation
$((a,0,\ldots 0);\;(b,0,\ldots 0);\;(c,0,\ldots 0))$
on the $\mathbb{Z}_{7}$-symbols read in positions $\vec{t}_{1},\ldots,\vec{t}_{M}$, where $\pi=(a\;b\;c)$.
Thus, $\psi_{0,n,\pi,0}$ is an element of the simple group of all even permutations of such vectors (performed on the $\mathbb{Z}_{7}$-components of cells $\vec{v}+T_{n}$ when $\vec{v}$ has a marker). For this to be well-defined, it is crucial that the automorphisms never change the positions of markers, and that the $\mathbb{Z}_{7}$-symbols being affected by the permutations cannot belong to markers at two different roots $\vec{v},\vec{v}^{\prime}$ such that both have a marker. But this is clear because we write the counters explicitly, never change the values of counters, and they describe the position of each cell in the corresponding marker $\vec{v}+T_{n}$, thus identifying the root $\vec{v}$.
We conclude that in $\mathrm{PAut}_{\mathbb{Z}^{D}}(A)$, there is an element of norm at most $C^{\prime\prime}n^{3}\log(n)^{t}$ for some global constants $C^{\prime\prime},t$, which belongs to a simple group $H$ of order $27^{M}!/$ 2 where $M=C^{\prime}n^{D-1}$ for a global constant $C^{\prime}$.
Suppose now for a contradiction that $\mathrm{PAut}_{\mathbb{Z}^{D}}(A)$ embeds in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for some alphabet $B$. By Lemma 4.1, there exists a constant $C$ such that if $H\leq\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ is a simple group containing a nontrivial element of norm at most $N=C^{\prime\prime}n^{3}\log(n)^{t}$, we have $|H|\leq C^{N^{d}}!$. But a short calculation shows that
$|H|=7^{C^{\prime}n^{D-1}}!/2>C^{(C^{\prime\prime}n^{3}\log(n)^{t})^{d}}!$
for large $n$ when $D-1>3d$. Namely, these are the respective degrees of double exponentials appearing in the formulas. This is a contradiction, and proves that no such embedding can exist.
To get the result for general alphabet $A^{\prime}\ni 0,1$, we can use words of the form $0110w$ 0 with $w$ not containing $11$ as a subword, to encode symbols of of $A^{2}$ where $A$ is the above alphabet (the square is for technical reasons, and the coding can be arbitrary), write $W$ for the set of these words, with $|W|=|A|^{2}$. Note that this set of words is mutually unbordered. The idea is that we use only the first encoded $A$-track in the argument, and the counters will be again written in consecutive positions in dimension $D$ as in the above proof, but now using $r(D+1)$ positions instead of $D+1$ as before. I.e. we use a constant number of such words $0110w$ 0 with $w\in\{0,1\}^{r-4}$ for fixed $r$ to encode the counters.
We replace the group $\mathrm{PAut}$ as follows: for symbol permutations, permute the symbols interpreted inside marker; as for partial shifts in direction $e_{i}$, we use the conveyor belt trick (see e.g. [28] for a more detailed discussion): if $i<D$, then in every chain of markers in direction $e_{i}$ (in the same phase, i.e. with the $11$-subwords in consecutive positions in direction $e_{i}\mathbb{Z}$), in each marker we interpret two symbols $s,s^{\prime}$ on the two tracks $A^{2}$, and shift the first symbol $s$ in direction $e_{i}$, and $s^{\prime}$ in direction $-e_{i}$. At the ends of such chains, we have a unique natural way to continue this to a bijection, by moving symbols between tracks, so they move as if on a conveyor belt. If $i=D$, we do the same, but now we look at progressions of consecutive markers with difference $e_{D}D$.
Now, the above argument goes through essentially verbatim, when we assume that all the controlled rotations are performed only on symbols on the first $A^{2}$-track (partial shifts may move $\mathbb{Z}_{7}$-symbols to the other track if they reach the end of the conveyor belts, but in this case, all the commutators will cancel in the end, and this $\mathbb{Z}_{7}$-symbol is not affected by the final automorphism). ∎
The case of groups containing free submonoids is similar, and we outline how to obtain it from the proof of Theorem 3.1.
Theorem 4.2.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet. Then $\mathrm{Aut}(A^{G})$ has an f.g. subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
Proof.
Assume $A$ has many prime divisors. In the case of a general alphabet, the marker argument is similar as in the previous theorem.
We explain only informally how the dyadically accelerated funneling argument is applied, i.e. the modifications made in the definition of the $\psi$-maps (and implicitly, the notion of goodness) in the proof of Theorem 3.1. First, the Turing machines $T_{1}$ and $T_{2}$ can now simply behave trivially, and accept all words. (These Turing machines play no role in this argument, and their states and tape symbols need not be included in the alphabets.)
Then $\psi$ simply rotates the $\mathbb{Z}_{7}$-symbol at the root of a tree where the nodes are all good and the other $\mathbb{Z}_{7}$-coordinates contain 00, and has polynomial norm in $n$. In particular, there is only one possible projection of such a tree on the non-$\mathbb{Z}_{7}$-coordinates, and furthermore, if we have such trees at $h,h^{\prime}$ with $h\neq h^{\prime}$, then $h\{a,b\}^{\leq n}\cap h^{\prime}\{a,b\}^{\leq n}=\emptyset$ (since the element of $\{a,b\}^{*}$ is visible in the index).
Thus, there is a well-defined group $H$ of even permutations on the $\mathbb{Z}_{7}$ track under such trees. Since the tree itself has $12^{n+1}-$ 1 nodes, the number of configurations is $7^{2^{n+1}-1}$, and thus we have a simple group of cardinality $|H|=7^{2^{n+1}-1}!/2$ with an element of polynomial norm. By Lemma 4.1, and a short calculation, we cannot have an embedding of $H$ into any group of the form $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$. ∎
Remark 4.1.
One could generalize results of this section to other subshifts and groups, with a bit of additional work, although we do not know whether this would lead to any striking results. There are some subtleties, e.g. in $\mathbb{Z}^{d}$-SFTs, it is surely possible that periodic points are dense, but patterns of shape $[0,n]^{D}$ do not fit into boxes of size computable in $n$. The proof of denseness of periodic points in FEP subshifts in [25] shows that at least on $\mathbb{Z}^{d}$, an analog of Lemma 4.1 does hold. In the case of general residually finite groups, even on full shifts it may not be possible to find efficient quotients to use in the proof. Furthermore, our proofs here also use rigid cones in the case of $\mathbb{Z}^{d}$ and embedded semigroups in the case of other groups. It should be straightforward to replace these arguments by the use of abstractly embedded trees, as in the following sections.
4.1 Dropping the centers
By completely analogous arguments, one can prove variants of the results where the center is quotiented out. We omit a technical version of Lemma 4.1, and explain the modifications needed directly.
The following is due to Hochman [19]. The one-dimensional result is due to Ryan [27].
Lemma 4.2.
The center of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ is the group of shifts.
Theorem 4.3.
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
Proof.
The proof is analogous to Theorem 4.1. Let $\mathcal{G}=\langle S\rangle$ be the group used in that proof. Then in $\mathcal{G}_{D}/K$ we take the the subgroup $\mathcal{G}/K$.
A we showed, there is a subgroup $H$ of $\mathcal{G}_{D}$ which is nonabelian simple of order $27^{\Theta(n^{D-1})}!/$ 2, and contains an element of word norm $n$. Now, $H$ must have trivial intersection with $K$, since otherwise $K\cap H$ is a nontrivial simple subgroup of $H$. Thus by the second isomorphism theorem, $HK/K\cong H$ as a subgroup of $\mathcal{G}/K$, and we conclude that $\mathcal{G}/K$ still has a simple subgroup of order $27^{\Theta(n^{D-1})}!/$ 2 and with an element of word norm $n$.
Next we observe that the set of groups embeddable in some $\mathcal{G}_{d}/K^{\prime}$ (over all alphabets $B$) does not decrease if we assume $K^{\prime}$ is just the subgroup $\langle e_{1},\ldots,e_{d^{\prime}}\rangle$ of $\mathbb{Z}^{d}$ for some $d^{\prime}\leq d$. Namely, we observe first (e.g. by linear algebra) that there is a subgroup $L$ such that $G=K^{\prime}\oplus L$ is of finite index in $\mathbb{Z}^{d}$. Any subgroup of $\mathcal{G}_{d}/K^{\prime}=\mathrm{Aut}(B^{\mathbb{Z}^{d}},\mathbb{Z}^{d})/K^{\prime}$ is certainly a subgroup of $\mathrm{Aut}(B^{\mathbb{Z}^{d}},G)/K^{\prime}$. Then we recall that the $G$-action $(B^{\mathbb{Z}^{d}},G)$ itself is topologically conjugate to a full shift on symbols $B^{[\mathbb{Z}^{d}:G]}$. By replacing the alphabet $B$ by this set and observing that $G\cong\mathbb{Z}^{d}$ (by the fundamental theorem of abelian groups), we may replace the original group $\mathbb{Z}^{d}$ by $G$. Since $K^{\prime}\times L$ is now an internal direct product in $\mathbb{Z}^{d}$, by the fundamental theorem of abelian groups $K^{\prime}\cong\mathbb{Z}^{d^{\prime}}$ and $K^{\prime}\cong\mathbb{Z}^{d^{\prime\prime}}$ for some $d^{\prime}+d^{\prime\prime}=d$, and thus by changing the basis we may assume $K^{\prime}$ is just the subgroup $\langle e_{1},\ldots,e_{d^{\prime}}\rangle$.
Assume now that $K^{\prime}$ is of this form, and let $L=\langle e_{d^{\prime}+1},\ldots,e_{d^{\prime\prime}}\rangle$. Consider a finitely-generated subgroup $H^{\prime}$ of $\mathcal{G}_{d}/K^{\prime}$. We may take representatives for the generators $\mathcal{S}^{\prime}\subset\mathcal{G}_{d}$, and suppose $r$ is a bound on their radii. Suppose that some nontrivial element $h\in H^{\prime}$ of this group has word norm $n$ and belongs to a nonabelian simple group of order $27^{\Theta(n^{D-1})}!/$ 2. Let $h^{\prime}$ be the corresponding representative in $\mathcal{G}_{d}$ (of the same word norm). Since $h$ is nontrivial, $h^{\prime}$ is not contained in $K^{\prime}$. Note that the radius of $h^{\prime}$ is at most $rn$.
We now claim that we can find a point with periods at most $10rn$ in each direction, such that $h^{\prime}$ does not act trivially on the $K^{\prime}$-orbits of periodic points. We observe that $\mathcal{G}_{d}/K^{\prime}$ has a well-defined action on the $K^{\prime}$ orbits of such points, and thus so does $H^{\prime}$. Since $H^{\prime}$ is simple, as in the proof of Lemma 4.1 we see that $H^{\prime}$ must act faithfully on the set of $K^{\prime}$-orbits, and we conclude that again $|H^{\prime}|\leq C^{n^{d}}!$ for some constant $C$ (which is only a function of $r$ and the alphabet. Note here that the constant $10$ also disappears into $C$, and that the number of $K^{\prime}$-orbits is smaller than the number of periodic points, which only gives a better upper bound for $H$.
There are two cases to consider for this claim. Suppose $B\ni 0,1$. First, suppose $h^{\prime}$ is a shift by an element not in $K^{\prime}$, then it is a shift by a vector of radius at most $rn$ from outside $K^{\prime}$. We can easily find point with periods $(2rn+1)$ where $h^{\prime}$ it does not act by a $K^{\prime}$-shift, for example we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}\in K$.
Suppose then that $h^{\prime}$ is not a shift map. Let first $y\in\{0,1\}^{\mathbb{Z}^{d}}$ be the characteristic function of $\{\vec{0}\}$, i.e. $y_{\vec{v}}=1\iff\vec{v}=\vec{0}$. Then we have two subcases. First, if $h^{\prime}(y)\notin K^{\prime}\cdot y$, then it is easy to see that we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}=\vec{0}$, and then $x$ is not mapped into its $K^{\prime}$-shift orbit.
Suppose then $h^{\prime}(y)\in K^{\prime}\cdot y$, so $h^{\prime}(y)=\sigma_{\vec{v}}(y)$ for some $\vec{v}\in B_{rn}\cap K^{\prime}$. Since $h^{\prime}$ is not a shift by an element of $K^{\prime}$, we can find a pattern $P\in B^{B_{rn}}$ such that $h^{\prime}(z)\neq\sigma_{\vec{v}}(z)$ when $z\in[P]$. Positioning $z|B_{rn}$ and a central pattern of $y$ far from each other in the same $10rn\mathbb{Z}^{d}$-periodic point $x$, we obtain a point which $h^{\prime}$ does not shift by an element of $K^{\prime}$.
This concludes the proof, since as explained above, this gives a bound of the form $|H|\leq C^{n^{d}}!$ as needed in the theorem. ∎
5 Arboreous graph families
In Section 6, we show that assuming the Gap Conjecture, nice enough SFTs on groups with superpolynomial growth have automorphism groups with PSPACE-hard word problem. The idea is to interpret trees (i.e. free monoids) in these SFTs. We cannot quite find full binary trees (this would be impossible in any group of subexponential growth), but from the growth rate assumption we can at least deduce the existence of polynomial-diameter trees that have a linear diameter binary tree as a path contraction (i.e. up to replacing paths with edges, we do have full binary trees).
In this section, we define the desired shapes of such trees in more detail, define an analog of $\mathrm{PAut}$ on graph families, and show that as soon as a graph family contains the right kind of trees, $\mathrm{PAut}$ groups can have PSPACE-hard word problems.
Definition 5.1.
Let $S$ be a finite set (think of it as the positive generating set of a free group). An $S$-labeled graph is a directed graph $G$ where each edge has a label from $S$ or $S^{-1}$. We say such a graph is good if
[…]
A graph is complete if for every vertex $x$, an outgoing edge exists with any label $s\in S^{\pm}$.
A weakly connected component of an $S$-labeled graph is just a connected component of the underlying undirected graph (with an undirected edge for each directed edge).
A good $S$-labeled graph can be thought of as a partial group action of the free group $F_{S}$. We also write $xs$ for the unique neighbor of $x$ where the edge has label $s$ (for $s\in S^{\pm}$), if this exists, and otherwise $xs$ is undefined.
Definition 5.2.
For $|S|\geq 2$, a depth-$k$ $S$-ary word tree is an $S$-labeled graph obtained as follows: Start with the graph with vertices $S^{\leq k}$ and an $s$-labeled edge $(u,us)$ for $s\in S,u\in S^{\leq k-1}$. Then for some $U\subset S^{\leq k}$, identify the vertices $\{uav,ubv,ucv\}$ whenever $u\in U,v\in S^{*}$, i.e. we quotient the tree by the minimal equivalence relation making these identifications. We say such a tree is an $(n,S)$-thicket if for each $u\in S^{k}$, there are at least $n$ prefixes $v\leq u$ such that |$|\{vs\;|\;s\in S\}|=|S|$ in the resulting tree.
We phrase the definition in this constructive way rather than descriptively, since otherwise it is difficult to state concisely that the tree “should not have any identifications other than ones arising from $ua=ub=uc$” (as certainly such an identification leads to many other identifications, but we only want to have the obvious consequences). One indeed obtains a full ternary tree from an $(n,\{a,b\})$-thicket by contracting the paths along vertices $u$ where $ua=ub$.
Definition 5.3.
Suppose $S\supset\{a,b,c\}$. A family of $S$-labeled graphs $\mathcal{G}$ is arboreous if there exist $C,d$ such that for all $n$, there exists $H\in\mathcal{G}$ and $h\in H$ such that the subgraph of $H$ induced by the vertices $h\{a,b,c\}^{\leq Cn^{d}}$ is an $(n,\{a,b,c\})$-thicket.
Next we define our $\mathrm{PAut}$ analog.
Definition 5.4.
Let $\mathcal{G}$ be a family of complete $S$-labeled graphs, let $F_{S}$ be the free group generated by $F_{S}$. For finite sets $A_{i}$, the group $\mathrm{PAut}_{\mathcal{G}}(A_{1},\ldots,A_{k})$ is defined by its action on $\bigsqcup_{G\in\mathcal{G}}A^{G}$ where $A=A_{1}\times\cdots\times A_{k}$. The generators are as follows:
  • For each $i\in[1,k]$ and $g\in F_{S}$, we have the partial shift $\sigma_{g,i}$ that maps $x=(x_{1},\ldots,x_{k})\in A^{G}$ to $(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{k})$, where $y_{h}=(x_{i})_{hg}$ for $h\in V(G)$.
  • For each $\pi\in\mathrm{Alt}(A)$, we have the symbol permutation $\eta_{\pi}$ that maps $x\in A^{G}$ to $y$ defined by $y_{h}=\pi(x_{h})$ for $h\in V(G)$.
We do not require that $\mathcal{G}$ has a unique copy of each graph up to isomorphism, but requiring this would not change the definition.
Theorem 5.1.
Let $\mathcal{G}$ be any arboreous path family. Then for some $A$, the group $\mathrm{PAut}_{\mathcal{G}}(A)$ has PSPACE-hard word problem.
Proof.
The proof is analogous to that of Theorem 3.1. We do not give it in detail, and only explain the main differences.
First, consider the case where $G$ contains a graph where for some vertex $h$, $h\{a,b,c\}^{*}$ is a copy of the ternary free monoid. In this case, we can directly use the arguments in the proof of Theorem 3.1, and perform dyadically accelerated funneling along the encoded ternary free monoids. For inputs $u$ where a witnessing tree exists (in the sense of Lemma A.3, proving $u\in L$), such a tree exists in some graph, and thus the $\mathrm{PAut}$-element is nontrivial.
If a witnessing tree does not exist, there is a slight subtlety, namely one may wonder if there can exist a graph where the tree is not injectively embedded, and this somehow leads to incorrectly thinking that it contains a witness. One can check that the logic of the proof does not assume that $hu\neq hv$ when $u,v\in\{a,b,c\}^{*}$ are two distinct words of the free monoid. The hard part is the inductive argument at the end, since there the same (first) track is acting both as control and as the permuted data. Here we note that the permutations are always commutatored with index checks, so the only situation where something nontrivial happens is that the index values are as they should along the computation tree. Thus, the controlling and permuted data is not mixed.
We conclude that the controlled permutation constructed in the end acts nontrivially precisely in the presence of a tree embedded on the graph.
Second, consider the general case, i.e. that we only have thickets with possibly polynomial paths between branchings in the trees. We still perform dyadically accelerated funneling, but now we need to think of the long non-branching paths as single edges, and find analogs of the automorphisms in the proof of Theorem 3.1 which can communicate over these paths.
The first thing we need is to have a way to check whether some node $h$ is branching, or more generally whether $hs=hs^{\prime}$ for $s,s^{\prime}\in F_{S}$. We can construct with the usual techniques first an automorphism that permutes the $\mathbb{Z}_{7}$-component at $h$ if the $\{0,1,2\}$-tracks at $hs$ and $hs^{\prime}$ have distinct values. Then we change the value at $hs$ and try again. One of these permutations succeeds if and only if $hs$ and $hs^{\prime}$ are distinct. The formulas from Section A.3 (or a variant of Equation 1) can now be used to obtain an automorphism that permutes the $\mathbb{Z}_{7}$-component at $h$ by a particular even permutation if and only if $hs\neq hs^{\prime}$.
Once we can check for equality of nodes, the proof is again similar to the proof of Theorem 3.1, but we check for an accepting tree on the tree obtained by contracting paths. We will construct versions of the $\phi$-maps that send information over paths where the tree is not branching, and goodness checks should similarly jump over paths. In the remainder of the proof, we explain how these variants can be constructed.
Note that the contracted paths are of length $n^{d}$ at most. We explain first how to deal with a single length $\ell$. A minor modification of Equation 1 gives an automorphism $f$ that performs $\pi$ in the $\mathbb{Z}_{7}$-component at $h$ if and only if $ha^{\ell}$ contains $s$. Namely, we simply replace the conjugating partial shifts of the third track by their $\ell$th power.
However, we need to check that the intermediate nodes on the path, i.e. $ha^{\ell^{\prime}}$ for $\ell^{\prime}<\ell$, do not branch, and that $ha^{\ell}$ does branch. This can be done since as we showed above, we have in our group automorphisms that permute $\mathbb{Z}_{7}$ in an arbitrary way depending on whether a given relative node is branching. Specifically, one can use commutator formulas and divide-and-conquer to obtain a polynomial-norm automorphism that cancels the effect of $f$ if the branching is not correct.
The resulting automorphisms can be simply composed for all distinct $\ell$ to obtain the desired analogs of the $\phi$-maps, as their supports are distinct (the $\ell$th map can only act nontrivially when the non-branching prefix of the tree starting from $h$ is of length exactly $\ell$).
To allow goodness checks to jump over paths, the argument is direct from Barrington’s theorem (taking the maps that check for branching nodes to be to be among the generators). ∎
6 PSPACE-hardness for general groups under the Gap Conjecture
A sequence $b_{k}$ is at most polynomial if $b_{k}=O(k^{d})$ for some $d$. A function $b_{k}$ is at least stretched exponential if we have $b_{k}=\Theta(e^{k^{\beta}})$ for some $\beta>0$, and $\beta$ is called the degree. We again recall the version of Gap Conjecture relevant to us, so let $\beta\in[0,1)$.
Conjecture 6.1 (Conjecture $C^{*}(\beta)$ of Grigorchuk).
A group either has at most polynomial growth, or at least stretched exponential growth with degree $\beta$.
We next show that in any group with stretched exponential growth in the above sense, for sufficiently large $k$ we can find a subtree that fits in a ball of radius polynomial in $k$, has sufficient branching, and furthermore we can encode this tree in a finite-support configuration of a particular SFT $X$.
We are mainly interested in full shifts, so we will work with SFTs that contain points of finite support, and our first result constructs finite-support points which we will use as markers. This does not require any assumptions on the group. Similar marker constructions can be performed under other (including some strictly weaker) assumptions. See e.g. [2, 19] for such marker constructions.
Lemma 6.1.
Let $X$ be a subshift of finite type where some finite-support point has free orbit. Let $C\in\mathbb{N}$ be arbitrary. Then we can find $k$ and a finite family $F$ of points with support contained in $B_{k}$, such that
  • There is a block map from $X$ to $\{0,1\}^{G}$ which maps every $x\in F$ to the characteristic function $\chi$ of $\{1_{G}\}$.
  • $|F|\geq C|B_{Ck}|^{C}$.
Proof.
Let $b_{k}=|B_{k}|$ and suppose $G$ has at most $s$ generators. Suppose $x\in X$ has support contained in an $m$-ball, and has free orbit, assume $m$ is a window size for $X$. It is reasonably easy to show that there is a block map $\phi$ which maps $x$ to the characteristic function $\chi$ of $\{1_{G}\}$. Namely, consider the block map whose local rule identifies the central $B_{m}$-pattern of $x$ by writing $1$, and otherwise writes 00. If some shift of $gx$ also has this pattern in $B_{m}$, then the nonzero symbols of $x|g^{-1}B_{m}$ are in bijection with those of $x$ in $B_{m}$. Thus actually $g^{-1}x=x$ because these must contain all the nonzero symbols.
We claim that there exists $\alpha>1$ such that for all $k$ we can find a family $F$ of points with support contained in $B_{k}$, and with cardinality at least $\alpha^{b_{k}}$, and such that some block map maps all of them to $\chi$.
To see this, first we may modify the local rule of $\phi$ so that it only outputs $1$ if it sees content from $x$ in the $10m$-ball. Thus, we can use the same block map, if we only use points where there is a copy of the central pattern of $x$ separated by at least $10m$ from other such copies, and these other copies are $3m$-dense in the annulus $B_{k}\setminus B_{10m}$ (which easily implies that each of them sees another one in the annulus $B_{k}\setminus B_{10m}$ for large enough $k$).
By the Packing Lemma we can indeed find a subset of the annulus which is $3m$-separated and $3m$-dense, and apart from the boundaries we can freely center copies of left translates of $m$-balls in these positions. Then we can encode information in the point by varying the exact positions by $1$ in various directions (we can move each pattern by at least $m/3$ before any two may intersect). In particular, we have a positive density of positions where we can write a bit of information, which easily implies the claim in the previous paragraph.
Now we claim that if $\alpha>1$, then $\alpha^{b_{k}}>Cb_{Ck}^{C}$ for some $k$, which is the claim about the size of $F$. Suppose not. Then for all $k$ we have
$\alpha^{b_{k}}<Cb_{Ck}^{C}.$
Taking logarithms we have
$b_{k}\log\alpha<\log C+C\log b_{Ck}\leq\log C+C^{2}k\log s$
so $b_{k}<(\log C+C^{2}k\log s)/\log\alpha$. Thus the growth of balls is bounded by a linear function of $k$. It is well-known that this implies that $G$ is virtually $\mathbb{Z}$, and in this case, the claim is clear. ∎
We now want to formalize the idea that we can interpret trees and other graphs on configurations of an SFT on a group, and furthermore on the nodes of these graphs we can interpret symbols that can freely be changed to others.
The following is a variant of [29, Definition 1]; in that definition, the class of graphs was specialized to be a family of cyclic groups, but additionally a subshift was allowed on these groups. A subshift could be allowed on the graphs here as well, but the definition is already quite complicated, and we only need full shifts in our application.
Definition 6.1.
Let $X\subset A^{G}$ be a subshift, $B$ a finite alphabet, and let $\mathcal{G}$ be a set of good $S$-labeled graphs. We say $X$ simulates $(B,\mathcal{G})$ if, possibly after applying a topological conjugacy to $X$, the following three paragraphs hold.
  1. 1.
[…]
  2. 2.
    To each $x\in X$ we can associate an $S$-labeled graph by taking nodes $V(X)=\{(g,j)\;|\;g\in G,0\leq j<m(x_{g})\}$, and for each $(g,j)\in V(x)$ with $v(x_{g})_{s,j}=(h,j^{\prime})$ adding an $s$-labeled edge from $(g,j)$ to $(gh,j^{\prime})$ (so we must have $(gh,j^{\prime})\in V(x)$). We require that this graph is good for all $x\in X$. Furthermore, the weakly connected components of these graphs are in $\mathcal{G}$.
  3. 3.
[…]
We say further that the simulation is complete if all the graphs are complete.
Lemma 6.2.
Let $G$ have at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT such that some finite-support point has free orbit. Then for any $B$ there exists an arboreous family of complete graphs $\mathcal{G}$ such that $X$ simulates $(B,\mathcal{G})$.
Proof.
Let $k$ be as in Lemma 6.1 for $C=\max(|B|^{2},15)$. Let $D=Cb_{Ck}^{C}$, and let $d$ be such that $\beta d>1$ where $\beta$ is the degree of stretched exponential growth. Then for some constant $c$,
$|B_{n^{2d}}|\geq ce^{(n^{2d})^{\beta}}=ce^{n^{2\beta d}}\geq(D+1)^{n^{2}}$
for large enough $n$ (since $n^{2\beta d}>2$).
Pick inside $B_{n^{2d}+k}$ disjoint balls $B^{i}=g_{i}B_{k}$, so that the set of centers $\{g_{i}\}$ is $3k$-dense and $3k$-separated. We may assume $g_{1}=1_{G}$ so that $B^{1}=B_{k}$. Connect each ball $B^{i}$ to all $B^{j}$ such that $d(g_{i},g_{j})\leq 7k$, by a two-directional edge. The resulting graph is connected and in fact the distance from $B^{1}$ to any other $B^{i}$ is at most $2n^{2d}$: we can follow a geodesic from $1_{G}$ to $g_{i}$ in the group $G$, and the sequence of $B^{j}$’s whose centers are closest to the geodesic (breaking ties arbitrarily) is connected by the edges of our new graph due to $3k$-denseness, and gives a path $B^{1}$ to $B^{i}$.
Since the packing is $3k$-dense, we have at least
$|B_{n^{2d}}|/|B_{3k}|\geq(D+1)^{n^{2}}/|B_{3k}|\geq D^{n^{2}}$
such centers for large $n$ (note that to fit balls inside $B_{n^{2d}+k}$, the centers should be in $B_{n^{2d}}$).
Next, take a spanning tree of our new graph, rooted at $B^{1}$, and with distance from $B^{1}$ to any $B^{i}$ at most $2n^{2d}$. For example, go through spheres of larger and larger radius, and arbitrarily connect (in the tree) every element of the $r$-sphere to an element of the previous sphere to which it is connected in the graph.
Next, associate to each node $B^{i}$ the length at most $2n^{2d}$ path from the root $B^{1}$ to it. We can code this path by the sequence of offsets between the centers of the balls $B^{j}$ on the path (as elements of $G$). This associates to $B^{i}$ a word of length at most $2n^{2d}$ over as set of size much smaller than $D$. We may then code this as a word over $\hat{D}=\{1,\dots,D\}$, using a tail over one of the symbols (say we use $D$ for this) to mark that the endpoint has been reached.
So at the moment, we have a set $L$ of words in $\hat{D}^{2n^{2d}}$, with $|L|\geq D^{n^{2}}$. Geometrically, this corresponds to a superexponentially large tree of polynomial depth, and we want to first conclude that such a tree must have a subtree where there are at least linearly many branchings along every path from the root to a leaf, and then from this to construct an $(n,\{a,b,c\})$-thicket.
One way to conclude the existence of subtrees that branch a lot is to use the winning set $\hat{L}$ of $L$, as defined in [30] (based on [33]). This is a set of words of the same length and over the same alphabet that can be associated to $L$, which is downward-closed, meaning if $v\in\hat{L}$ and $u\leq v$ in the sense that $u_{i}\leq v_{i}$ for all $i\in[0,2n^{2d}-1]$, then $u\in\hat{L}$.
The winning set $\hat{L}$ encodes “branching structures of uniformly branching trees” in $L$, meaning that if $v\in\hat{L}$, then letting $U=\{u\;|\;u\leq v\}$, there is an assignment $T_{v}:U\to L$ such that for all $u,u^{\prime}\in U$, we have
$\min\{i\;|\;u_{i}\neq u^{\prime}_{i}\}=\min\{i\;|\;T_{v}(u)_{i}\neq T_{v}(u^{\prime})_{i}\}.$
Thus, geometrically, in the spanning tree $v\in\hat{L}$ corresponds to a subtree where whenever $v_{i}\geq 2$, the tree branches in at least two in each branch at depth $i$.
Now the claim about tree branching linearly many times on each path amounts to showing that we can find $v\in\hat{L}$ with support size at least $n$. We show that indeed this is immediate from the cardinality of $\hat{L}$ (which is that of $|L|$), by some simple combinatorics.
Let $s$ be maximal number of non-$1$ coordinates in a word $v\in\hat{L}$. Then to describe words in $\hat{L}$ it suffices to list the $s$ positions of non-$1$ symbols (or a superset of them) and describe their contents. Thus we have
$D^{n^{2}}\leq|L|=|\hat{L}|\leq D^{s}\binom{2n^{2d}}{s}\leq D^{s}(2n^{2d}e/s)^{s}$
from Stirling’s approximation. Taking logarithms we have
$n^{2}\log D\leq s(\log D+\log(2n^{2d}e/s))$
Since $\log(2n^{2d}e/s)$ grows much slower than $\sqrt{n}$, we even have $s=\Omega(n^{1.5})$ for large $n$, meaning we actually have much more than a linear amount of branching in some subtree.
To get an $(n,\{a,b,c\})$-thicket, we actually need the brachings to be ternary, i.e. along every path to a leaf, we should have many nodes with three children. This is not automatic, but by using paths of length at most $2$ of the tree (call these $2$-jumps) instead of the original edges (which correspond to length-$14k$ elements of $G$), it is possible to mimic the usual proof that $\{a,b,c\}^{*}$ embeds in $\{a,b\}^{*}$ as follows. First, as some preprocessing, step into a subtree so that every branching is binary, and drop every second branching so that there is a path of length at least $2$ between any consecutive branchings.
In the resulting tree, go down the tree from the root, and at the first branching (let us say to a “left” and a “right” child), use the left branch as such with label $a$, but in the right branch, have an edge with label $b$ to the original right neighbor, and a $2$-jump with label $c$ to the unique child down the tree from the right neighbor. This gives the node three successors connected by $2$-jumps. In the left branch, proceed to modify the tree recursively. In the right branch, continue down the original path with $2$-jumps to simulate two paths on one actual path. At the next branching, the two paths can merge into the successors, and we can recursively continue on these two paths. This again at most halves the number of branchings in the tree, and now every branching is ternary.
Since $s=\Omega(n^{1.5})$, certainly $s/4\geq n$, so we have shown the existence of an $(n,\{a,b,c\})$-thicket, the labels being given as in the previous paragraph. Note that now the branchings are no longer uniform (unlike in the tree given by the winning set argument), but this is not required in the definition of a thicket.
By the choice of $k$, we can now position finite $k$-support patterns, namely central patterns of the set $F$ from Lemma 6.1, in the balls $B^{i}$ of the tree, where $|F|\geq C|B_{Ck}|^{C}$ for $C=\max(15,|B|)$. Distances in the tree of $B^{i}$s correspond to elements of $G$ with word norm at most $7k$, so double jumps correspond to elements with word norm at most $14k$, so it suffices to code four elements of $B_{14k}$ in each $B^{i}$, for which $|F|\geq C|B_{Ck}|^{C}$ is plenty since $C\geq 15\geq\max(4,14)$. We also still have room to store two symbols from the alphabet $B$, which is possible since $C\geq|B|^{2}$. Recall that a local rule can detect the positions of these patterns, and they do not overlap (by $3k$-separation), so automorphisms can freely change the $B$-symbol without messing up the encodings.
To fit this more explicitly into Definition 6.1 above, we can apply a topological conjugacy to $X$ which, when a central pattern of one of the points $F$ is visible (meaning the local rule $\phi$ outputs $1$) and at distance at least $3k$ from other such patterns, writes down the index of this point as a number between $\{1,\ldots,|F|\}$, which we assume disjoint from the alphabet $A$ of $X$, and maps the rest of the ball to a new symbol ##\##.
Clearly writing back the $F$-patterns gives a left inverse for this block map, so the map is injective, and thus the image is an SFT $Y$. In this $Y$, some positions of some configurations have a number from $\{1,\ldots,|F|\}$ that can be freely changed. These are the positions where we will take $m(c)\geq 1$ in the simulated graph.
To get not necessarily complete graphs, we could use $m(c)=1$ always. However, we want the graphs to be complete, while not removing any subgraphs, thus not removing the $3$-thickets whose existence we showed.
For this we again can use the conveyor belt trick: we take instead $m(c)=2$ always (so there are two simulated nodes in each $B^{i}$), and we wrap incomplete $s$-labeled paths at the ends into a conveyor belt. This is indeed why we store two symbols in each node instead of one. Note also that in the conveyor belt construction, we do not need to store more edges, as the two simulated nodes use the same edges, only flipping which edges is incoming and which is outgoing.
Indeed the $3$-thickets still appear on the encoded graphs, on the first simulated nodes where edges go forward. ∎
Combining the above lemma with Theorem 5.1 gives the desired theorem.
Theorem 6.1.
Let $G$ have at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT such that some finite-support point has free orbit. Then $\mathrm{Aut}(X)$ has PSPACE-hard word problem.
Corollary 6.1.
Let $G$ be a finitely-generated group with at least stretched exponential growth. Then $\mathrm{Aut}(A^{G})$ has PSPACE-hard word problem whenever $A\geq 2$.
7 The lamplighter group
Let $C$ be a finite nontrivial abelian group. The lamplighter group is $L=C\wr\mathbb{Z}=\bigoplus_{\mathbb{Z}}C\rtimes\mathbb{Z}$, where in the latter representation, $\mathbb{Z}$ acts by shifting the coordinates in $\bigoplus_{\mathbb{Z}}C$.
Theorem 7.1.
Let $C$ be a finite nontrivial abelian group, let $A$ be a nontrivial finite alphabet, $n\geq 2$, and $L=C\wr\mathbb{Z}$. Then $\mathrm{Aut}(A^{G})$ has co-NEXPTIME-complete word problems.
Proof.
We will assume $C=\mathbb{Z}_{2}$, because all the cases are similar (or rather, $\mathbb{Z}_{2}$ is the hardest case, as it needs a preprocessing step due to insufficient branching). We will also assume that $A$ is of specific form, as in the general case one can argue as in the proof of Theorem 4.1, i.e. work with a set of unbordered words representing symbols from a larger alphabet.
[…]
Now as the alphabet we take $A=\mathbb{Z}_{5}\times\mathbb{Z}_{6}\times\mathbb{Z}_{7}\times W\times\{0,1,2\}^{2}\times B\times D$, where $B=C^{2}$ and $C$ is the set of colors of Wang tiles, and $D$ is an auxiliary finite set.
The basis of the proof is the same as in [4]. The lamplighter group contains large grids. More specifically, geometrically this group is a horocyclic product of two full binary trees, which means that we can obtain it as a connected component of a direct product of such trees, where one tree is inverted, and we have to move up and down simultaneously in the two trees. (Thus when in one tree a node branches, in the other we move to a parent.) Now, a full binary tree of depth $n$ naturally has $2^{n}$ elements on its frontier, and we can identify them with the numbers $[0,2^{n}-1]$. In the direct product, there is then a natural copy of a square $[0,2^{n}-1]^{2}$. The square happens to be in a single connected component of the horocyclic product, which will allow us to use automorphisms to ensure computation on the massive grid obtained by connecting ($\mathbb{Z}^{2}$-)neighboring elements of the square.
It is useful to step into a subgroup before describing the grids formally. Let $G$ be an index-$2$ subgroup of $L$ isomorphic to $\mathbb{Z}^{2}_{2}\wr\mathbb{Z}$. Write $L^{\prime}$ for the natural copy $\mathbb{Z}_{2}\times\{0\}\wr\mathbb{Z}$ of $L$ in $G$. Let $t$ be the natural generator of $\mathbb{Z}$, and write $a=t$, $b=t\cdot(1,0)$, $c=t\cdot(0,1)$ where $(1,0),(0,1)$ are the natural generators of $\mathbb{Z}_{2}^{2}$. So $a,b$ are a standard set of generators of $L^{\prime}$. The point of this is that for each $g\in L^{\prime}$, $g+\langle c\rangle$ is a copy (coset) of $\mathbb{Z}$ where we can encode counters, and for distinct $g,g^{\prime}\in L^{\prime}$ these cosets are disjoint.
Let $g\in L$. The grid (of size $2^{n}$ at $g$) is the function $f:\mathbb{Z}_{2^{n}}\times\mathbb{Z}_{2^{n}}\to gL^{\prime}$, which for $u=\mathrm{bin}(\ell),v=\mathrm{bin}(m)$ maps
$f_{g}(\ell,m)=g\prod_{\{i\;|\;u_{i}=1\}}a^{-i}ba^{i-1}\prod_{\{j\;|\;v_{j}=1\}}a^{j}b^{-1}a^{-j+1}.$
Write also $f(\ell,m)=f_{1}(\ell,m)$ so $f_{g}(\ell,m)=gf(\ell,m)$. Here, words are $1$-indexed and recall that binary representations are written with least significant bit first, i.e. LSB. Sometimes, we call the image of this function the grid, and $g$ is the root of the grid.
Note that $f(\ell,m)$ for different $\ell,m$ commute as elements of $L^{\prime}$ (the total shift of such an element is 00, so they belong in the $\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}$-part of $\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}\rtimes\mathbb{Z}$), and form a copy of $\mathbb{Z}_{2}^{(2^{n})^{2}}$ inside it, whose action can be described (through the identifying map $f$) as swapping adjacent rectangles at different levels. We think of the grid as being drawn on a plane $\mathbb{R}^{2}$ with north, south, west and east as the four directions. The elements $a,b$ move “up” in this point of view, For a picture of a piece of this group (drawn as described here), and a highlighted grid, see [4].
The counters of $g\in L$ are the two numbers read in the two first components of the $\{0,1,2\}^{3}$-component of the alphabet, read in positions $g,gc,\ldots,gc^{n-1}$, and interpreted in binary (if there are $2$s present, the counters are undefined).
The basic idea (which we will have to modify a bit) is to write an element of $\mathrm{PAut}(A)$ of polynomial word norm in $n$, which works as follows: For $g\in G$, it performs a nontrivial $\mathbb{Z}_{7}$-permutation if and only if on the grid of size $2^{n}$ at $g$,
  • there is a valid tiling of $W$ in the $W$-components of the alphabet;
  • the counters at $f_{g}(\ell,m)$ have value $\ell,m$ respectively;
  • on the southmost row, i.e. in $f_{g}(0,0),f_{g}(1,0),\ldots f_{g}(2^{n}-1,0)$, we have a specific initial configuration (the initial configuration we are reducing to the word problem of $\mathrm{PAut}(A)$) for the Turing machine being simulated by $W$ (with polynomially many initial cells having specified content, and then only padding);
  • on the northmost row we have a configuration representing an accepting computation of $T$.
If we can write a formula for such an element in polynomial time, then we have a reduction from the problem of checking whether a length $2^{n}$-computation of $T$ accepts the given input of size $n$, to the co-word problem of $\mathrm{Aut}(A^{L})$. The problem being reduced is indeed NEXPTIME-complete since $T$ is a universal Turing machine, and because Wang tiles simulate the Turing machine explicitly, i.e. they directly code a spacetime diagram of the Turing machine with the (south-north) height of the tiling corresponding exactly to computation time, and (west-east) width exactly to space usage. Note also that as in the case of the PSPACE-hardness proof in Lemma A.3, while we allow only an exponential time rather than $2^{O(n^{d})}$ as in the definition of NEXPTIME, we can always produce a polynomial padding in the reduction.
As in previous hardness proofs, it is difficult to an element as described above directly, since the grid is not directly connected inside $L$, and we have to do all the checking outside the grid. Again, we will use dyadically accelerated funneling.
More specifically, observe that for fixed $i\in[0,2^{n}-1]$ we have
$f([0,2^{n}-1],i)=f(0,i)a^{-n}\{a,b\}^{n}.$
Thus, to connect $f(0,i)$ to the entire row $f([0,2^{n}-1],i)$, we can move by a right translation by $a^{-n}$, and then simply traverse up the full binary tree $\{a,b\}^{n}$. Note that in our geometric setting, this connecting takes place below the grid.
Similarly, there is a tree above the grid connecting each column, but we will not give the formulas and argue instead by symmetry, noting that the lamplighter group admits the automorphism $a\mapsto a^{-1},b\mapsto b^{-1}$ that inverts the point of view.
Just like in the proof of Theorem 3.1 we read words over $(0,1)^{*}2^{*}$ (indices) written on $(hc^{0},hc^{1},\ldots,hc^{n-1})$ for $h\in\{a,b\}^{\leq n}$ in the first $\{0,1,2\}$-component, and construct an automorphism $\psi_{1}^{\prime}$ of polynomial norm in $n$ that performs a nontrivial permutation $\pi$ at position $g$ in the $\mathbb{Z}_{7}$-component, if the positions $g\{a,b\}^{n}$ encode a valid (west-east row of a) $W$ tiling.
The basic idea is the same, i.e. an inductive argument using dyadic intervals $I$ of depths, which are read on the indices. However, now instead of having Turing machine configurations rooted at the various $gh$ with $h\in\{a,b\}^{\leq n}$, we have in each node just a single $B$-symbol, which codes $C^{2}$ where $C$ is the set of colors of $W$. When going up the tree, from $(c,c^{\prime})$ we produce $(c,c^{\prime\prime})$ in the $a$-branch and $(c^{\prime\prime},c)$ in the $b$-branch. The interpretation of $(c,c^{\prime})$ in a node is that $c$ (resp. $c^{\prime}$) is the west (resp. east) color of the leftmost leaf eventually produced from that node. The effect of this is to make the west and east colors consistent on the leaves. On the final $n$th level (which is where the grid of interest is), we check that the colors $(c,c^{\prime})$ are the left and right color of the Wang tile present in that node (it does not matter what the $W$-components contain outside the grid).
We should also ensure at this point that the final row has an accepting configuration. This is easy to include this in the check in the topmost dyadic interval of size $2$.
In addition to synchronizing the colors, the indices are directly used to synchronize the first $\{0,1,2\}$-components of the counters (to the correct grid value, relative to $g$). In the second $\{0,1,2\}$, we check that the tree contains a constant counter value, so that the south-north coordinate stays fixed.
The automorphism $\psi_{1}^{\prime}$ now performs some permutation ($\pi$) $n$ steps below valid grid rows whenever the grid coordinates and Wang tiles are consistent on that row, and the tree below has suitable content. We want to “move” these permutations to the left border on the grid, i.e. by $a^{n}$. We will move them to the $\mathbb{Z}_{6}$-component: use a partial shift of the $\mathbb{Z}_{7}$-track to move the (possibly) permuted cell of $\psi_{1}^{\prime}$ by $a^{n}$, conditionally permute the $\mathbb{Z}_{6}$-track, and shift back. Doing this twice, analogously to Equation 1, we can detect nontrivial permutations of $\mathbb{Z}_{7}$ and convert them to a nontrivial permutation of $\mathbb{Z}_{6}$ under the same conditions (but now happening on the grid).
Note that for this to work, it is important that the $\mathbb{Z}_{6}$-content is not read or affected by $\psi^{\prime}$ (which is the case); and that the $\mathbb{Z}_{7}$-cells on a valid grid are never affected by $\psi_{1}^{\prime}$. The latter holds because when $\psi_{1}^{\prime}$ acts nontrivially, counters on the grid contain binary numbers in the first counter, while any cell permuted by $\psi_{1}^{\prime}$ has has $2$ in its first $\{0,1,2\}$-component (being the root of a connecting tree, thus having index $2^{n}$ in the terminology of the proof of Theorem 3.1).
From this we obtain an automorphism $\psi_{1}$ that performs a nontrivial permutation on the $\mathbb{Z}_{6}$-component at $g$ if the tiles $f_{g}([0,2^{n}-1],0)$ match (west-east) horizontally, the counter values are consistent in the sense that the value at $f_{g}(i,0)$ is $(i,j)$ (and $j$ stays constant on the row), the nodes on the tree below have correct indices, and $\mathbb{Z}_{7}$-data (all zero, except the root having some element from the support of $\pi$).
By symmetrically using trees that connect the grid from “above”, we have elements $\psi_{2}$ that only act nontrivially at $g$, in the $\mathbb{Z}_{6}$-component, if the “column” $f_{g}(0,j)$ has consistent colors in the Wang tiles, in the south and north directions, and the counters are also are consistent.
In the argument giving $\psi_{1}$, the basis is that we can perform a controlled permutation on the $\mathbb{Z}_{7}$-component based on coordinates being good (though here we only implicitly defined the notion of goodness). These formulas are ultimately commutator formulas, and the goodness checks are ultimately of the form “perform a permutation $\pi^{\prime}$ if the node is good”. Thus, we can equally well use the map $\psi_{1}$ in place of the notion of goodness, but now using the $\mathbb{Z}_{6}$-components, and use the trees above the grid connecting columns from above, analogously to the definition of $\psi_{2}$. Then we can use partial shifts (as in the transition from $\psi_{1}^{\prime}$ to $\psi_{1}$) to move this permutation to the $\mathbb{Z}_{5}$-component of a node on the grid the grid. This gives us an automorphism $\psi_{3}$ that performs a nontrivial permutation in the $\mathbb{Z}_{5}$ component at $g$ if and only if every west-east row of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ rooted at $g$ is consistent (individually). Symmetrically from $\psi_{2}$ we can build, analogously to the definition of $\psi_{1}$, an automorphism $\psi_{4}$ that acts nontrivially if and only if all the columns are valid.
Now the commutator $[\psi_{3},\psi_{4}]$ has word norm $n$ in $\mathrm{PAut}(A)$, and performs a nontrivial permutation at $g$ if and only if every west-east row and every south-north column of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ has consistent tiling and consistent Wang tiles.
Finally, by taking another commutator, we can additionally condition the permutation by the initial configuration. Using the tile $\square$, we can in NC1 force the exponentially long tail to be $\square$ after the input.
Thus, we have reduced the NEXPTIME-complete problem of existence of $2^{n}$-time $T$-computations on polynomial-sized inputs (through the existence of tilings) to the complement of the word problem of $\mathrm{PAut}(A)$. ∎
8 Upper bounds and polynomial splitting schemes
In this section we show the co-NP, PSPACE and co-NEXPTIME upper bounds. In the case of co-NP, we also show completeness in this section (as this requires only a single sentence).
8.1 The co-NP upper bound
Theorem 8.1.
For an f.g. group $G$ with polynomial growth, the automorphism group of an SFT with NP language has co-NP word problem.
Proof.
Polynomial growth is equivalent to virtually nilpotent, by Gromov’s theorem. It is known that these groups have polynomial-time word problem. The word problem of any f.g. subgroup can be solved by guessing a witness (which fits in a polynomial-sized ball due to the growth rate). Checking the witness is in the SFT is in NP by assumption, and checking that the automorphism acts nontrivially on it can be done in polynomial time by applying the local rules. ∎
Corollary 8.1.
For an f.g. group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ always has co-NP-complete word problem.
Proof.
By the previous theorem, the word problem is in co-NP, and since virtually nilpotent groups contain copies of $\mathbb{Z}$, the hardness result for $\mathbb{Z}$ shown in [29] implies the same for $G$. ∎
8.2 The PSPACE upper bound
We next state a theorem that implies PSPACE upper bounds for word problems of automorphism groups of full shifts in the case of free groups and surface groups. The idea is that in these groups, when looking for a witness for the co-word problem (i.e. a pattern where the given element acts nontrivially), we can split the ball of the group in two pieces of roughly equal size, and continue recursively on the two halves. In the case of a free group, we can use the fact that the Cayley graph is a tree (so we can use constant-size cut sets), and in the case of surface groups, we can use geodesics to obtain linear cuts. To make the idea precise and to (perhaps) allow applying it to other groups, we give a general definition of polynomial splitting schemes.
Let $S$ be a fixed symmetric generating set containing $1_{G}$ for the finitely-generated group $G$.
Definition 8.1.
A set $A\subset G$ admits an $\alpha$-splitting scheme by $k$-cuts if $A=\emptyset$, or there is a set $C$ of cardinality at most $k$ such that $A\setminus C=A_{1}\cup A_{2}$ such that $A_{1}S\cap A_{2}S=\emptyset$, $|A_{1}|\leq\alpha|A|$, $|A_{2}|\leq\alpha|A|$ and each $A_{i}$ admits an $\alpha$-splitting scheme by $k$-cuts. We say $G$ admits polynomial splitting schemes if there exists $d$ and $\alpha<1$ such that $B_{n}$ admits an $\alpha$-splitting scheme by $(n^{d}+d)$-cuts for all $n$.
It is easy to see that this does not depend on the generating set (by thickening the splitting sets defined with respect to one generator), more generally see Lemma 8.3.
Theorem 8.2.
Let $G$ be an f.g. group with PSPACE word problem. If $G$ admits a polynomial splitting scheme, the automorphism group of any PEPB subshift has PSPACE word problem.
Proof.
We first consider a full shift $A^{G}$, and explain how to extend this to FEP subshifts. It suffices to show that the co-word problem is in alternating polynomial time AP, since co-PSPACE = PSPACE = AP (by determinism and Lemma A.2). Let $F$ be a finite set of automorphisms, and let $r$ be a bound on their radii. We may assume $r=1$ by changing the generating set $S$ of $G$. Given a word $w$ of length $k$ over $F$, we now describe how to prove the nontriviality, equivalently that there exists a pattern $p\in A^{B_{k}}$ such that composing the local rules of the cellular automata $w$ in sequence and applying them to $p$ gives a symbol distinct from $p_{1_{G}}$ at $1_{G}$.
The witness can be taken to be a pattern in $A^{B_{k}\times[0,k]}$ such that the patterns $c_{i}=(g\mapsto A^{(g,i)})\in A^{B_{k}}$ are obtained by successively applying the generators in the word $w$ to $c_{0}$ (and the symbols on the border not given by the local rule are filled arbitrarily). The contents of $B_{k-i}$ in $c_{i}$ are meaningful, while the rest is more or less meaningless, but makes indexing a little easier. At the end we should check that the contents of $1_{G}$ in $c_{k}$ differ from the contents of $1_{G}$ in $c_{0}$.
We show that we can figure out in alternating polynomial time whether such a pattern exists. To do this, we guess a polynomial sized subset $C_{1}$ of $B_{k}$. Then we compute its (polynomial-sized) boundary $C_{1}S\setminus C_{1}$. Note that we can check in polynomial space, and thus alternating polynomial time, whether two elements of the boundary are connected, and thus we can figure out a polynomial-size set of representatives of the boundary (or we can use $C_{1}S\setminus C_{1}$ directly and redo much of the computation). [^2] There is a little subtlety here: In alternating polynomial time computations, we may use PSPACE computations in recursive calls, which means we have alternating access to PSPACE. It may seem that this implies that APTIME trivially contains APSPACE, contradicting the fact APSPACE = EXPTIME and EXPTIME ${=}$ PSPACE = APTIME is a major open problem. The point is that alternating access from a polynomial time computation to PSPACE does not mean we have access to APSPACE, as in the latter we can do exponentially long computation including alternations. We also guess a pattern $p\in A^{C_{1}\times[0,k]}$ at this point.
The point is that we want to now check for each boundary element $g$, whether, letting $D$ be the connected component of the Cayley graph of $B_{k}\setminus C$ containing $g$, we can find a pattern $q\in A^{D\times[0,k]}$ which agrees with $p$ in $C\times[0,k]$, everywhere follows the local rules (in the sense described two paragraphs above), and that the contents of $1_{G}$ in $c_{k}$ differ from the contents of $1_{G}$ in $c_{0}$ if $1_{G}\in D$. The algorithm may not be correct for all $C,g$, [^3] In case we are not following a good sequence of splits, we may not shrink the set quickly enough, and the algorithm itself cannot easily check how much it shrinks the set at each point. but its positive claims are always going to be correct.
To do this, we recursively check for each component $D$ of $B_{k}\setminus C_{1}$ (represented by $C_{1}$ and an element of $C_{1}S\setminus C_{1}$) that we can find a polynomial-sized cut $C_{2}$ of $D$ which is contained in that component (again, use polynomially many PSPACE computations to ensure that $C_{2}$ is indeed contained in $D$, by finding paths to the boundary element $g$). When stepping forward to a connected component, we guess an extension of the previous pattern $p$ to $C_{2}$, and check that the local rules are followed everywhere.
Continuing this way, if a polynomial splitting scheme exists, in about $\log_{\alpha}|B_{k}|$ steps of recursion (which amounts to a linear amount of steps in $k$ since $\alpha>1$) we have figured out whether a witness exists: If one exists, we can use existential steps to ensure that the $C_{i}$ follow the polynomial splitting scheme, and that the patterns guessed on the $C_{i}$ come correspond to a pattern where $w$ acts nontrivially; when stepping to connected components, we use universal steps to ensure that all of the components admit such valid contents. On the other hand, the algorithm can only succeed if it manages to both find a polynomial splitting scheme, and a valid configuration where $w$ acts nontrivially.
To extend the result to PEPB subshifts, we find a locally valid pattern on $B_{n^{t}}$ where $t$ is the PEPB degree, and the algorithm in all branches checks local validity of the pattern $c^{0}$. Note that $((n^{t})^{d}+d)$ is still polynomial in $n$, so the polynomial splitting scheme for the $n^{t}$-ball still has cuts of polynomial size. By the definition of PEPB, the central $B_{k}$-part is globally valid, so the algorithm correctly identifies valid patterns where automorphisms act nontrivially. ∎
The main examples we have of groups with polynomial splitting schemes are free groups and surface groups.
Lemma 8.1.
Free groups have polynomial splitting schemes.
Proof.
For a free group, the Cayley graph is a tree, and as the cuts one can use successive elements of geodesics: at each step, we have a single branch of the ball rooted on some $g$, and stepping forward by one step in the unique geodesic from $1_{G}$ to $g$ splits the branch in three sets of equal size, which gives a polynomial splitting scheme for $\alpha=2/3$. ∎
In the case of surface groups, we assume that the reader is familiar with some basic notions from geometric group theory. A surface group is the fundamental group of a closed surface with some genus $g$. In the case $g=1$, the group is $\mathbb{Z}^{2}$ and word problems in automorphism groups are in co-NP. The surface groups with $g\geq 2$ are known to be commensurable, and nonamenable. For $g=0$ the group is trivial and $g<0$ reduces to $g>0$ after taking a finite cover (giving a commensurable group).
Definition 8.2.
The pentagon model of the hyperbolic plane is: the graph with vertices $\mathbb{Z}^{2}$, and undirected edges $\{(x,y),(x+1,y)\}$, $\{(x,y),(2x,y+1)\},\{(x,y),(2x+1,y+1)\}$.
Lemma 8.2.
All nonamenable surface groups are quasi-isometric to the pentagon model.
Proof.
As discussed above, it suffices to consider $g\geq 2$, since nonamenability is preserved by commensurability. Nonamenable surface groups are quasi-isometric to the hyperbolic plane [11]. Specifically, it is classical that the universal cover of the surface is the hyperbolic plane $H^{2}$, and by Švarc-Milnor the fundamental group is quasi-isometric to it. On the other hand, the hyperbolic plane is well-known to be quasi-isometric to the pentagon model. ∎
It should be clear how to generalize polynomial splitting schemes to any undirected graph, using the standard notion of cutsets (of vertices) in graphs.
Lemma 8.3.
The existence of polynomial splitting schemes in a graph is invariant under quasi-isometry.
Proof.
Balls of one graph are sandwiched between balls of the other graph by the quasi-isometry, so the quasi-isometry maps polynomial splitting schemes of a ball to polynomial splitting schemes of some shape that contains a large ball of some smaller radius. By restricting it to the ball, we obtain a polynomial splitting scheme for it. The radii change only by multiplication with the QI constants, so being of polynomial size retains its meaning. ∎
Lemma 8.4.
Surface groups have polynomial splitting schemes.
Proof.
It suffices to show that these exist in the pentagon model. Furthermore, it is easy to reduce the problem to finding a polynomial splitting scheme for “one-sided balls” $P_{k}=\{(x,y)\;|\;y\in[0,k],x\in[0,2^{y}-1]\}$, i.e. vertices reachable from $(0,0)$ by only moves that increment the second coordinate. Namely, these sets sandwich contain balls of the pentagon model, of roughly the same radius.
For the set $P_{k}$, a polynomial splitting scheme is obtained from geodesics between $(0,0)$ and vertices of the form $(x,k)$ with $x\in[0,2^{y}-1]$. Specifically, identify $(x,k)$ with $x/2^{k}$ and repeatedly cut the real interval in half. ∎
Lemma 8.5.
If $G$ has a polynomial splitting scheme, so does $G\times H$ whenever $H$ has polynomial growth.
Proof.
Taking $S\times T$ as a generating set (which by Lemma 8.3 does not affect the existence of polynomial splitting schemes), a ball of $G\times H$ is a product of balls of $G$ and $H$. Balls $B_{k}^{H}$ are polynomial size in $H$, so the cutsets $C$ for balls of $G$ appearing in the definition of a polynomial splitting schemes can simply be replaced by $C\times B_{k}^{H}$. ∎
Theorem 8.3.
For nonamenable free groups, nonamenable surface groups and their products with virtually nilpotent groups, the automorphism group of any PEPB subshift has PSPACE-complete word problem.
Proof.
These groups have polynomial splitting schemes, so the word problem of the automorphism group of any PEPB subshift is in PSPACE by Theorem 8.2. By Theorem 3.1, for the f.g. subgroup $\mathrm{PAut}$ the problem is PSPACE-hard in these cases, since nonamenable free groups and surface groups have free submonoids. ∎
8.3 The co-NEXPTIME upper bound
We now give the upper bound for the word problem that “always” holds, i.e. co-NEXPTIME.
Theorem 8.4.
Let $G$ be an f.g. group. The automorphism group of any subshift $X\subset A^{G}$ has co-NEXPTIME word problem under either of the following conditions:
  • $G$ has co-NEXPTIME word problem and $X$ has PEPB;
  • $G$ has EXPTIME word problem and $X$ has NP language.
In particular, if $G$ has word problem in P (as natural groups tend to do), then it has EXPTIME word problem, but EXPTIME is strictly larger than P.
Proof.
We start with the first item, i.e. assume $G$ has co-NEXPTIME word problem and $X$ has PEPB with degree $t$. Let $S$ be the generating set of $G$. We show that co-word problem (nontriviality of a given element) is in NEXPTIME. Given $w$ over a generating set of automorphisms with radius $r$, guess a good $S$-labeled directed graph $H$ (which is supposed to be $B_{(nr)^{t}}$) which is equal to the $(nr)^{t}$-ball around some $h\in H$ in $H$.
For every pair of distinct elements $h,h^{\prime}\in B_{(nr)^{t}}$, check that $h^{-1}h^{\prime}$ is nontrivial in $G$. This check can be done in NEXPTIME for each individual pair (it takes at most $e^{O(((nr)^{t})^{d})}=e^{O(n^{td})}$ steps for some $d$), and since there are at most $|S|^{{nr}^{t}}$ pairs it can be done in $|S|^{(nr)^{t}}e^{O(n^{td})}=e^{O(n^{t(d+1)})}$ time in total. If the check succeeds, we do not necessarily know that $H$ is equal to the actual ball $B_{(nr)^{t}}$, but we can be sure that $B_{(nr)^{t}}$ has the graph $H$ as an image under an edge label preserving graph homomorphism (or in more group-theoretic terms, $B_{(nr)^{t}}$ covers $H$).
Now guess a pattern $p\in A^{H}$. Check that it has no forbidden patterns (i.e. that it is locally valid), and that applying the local rules of the automorphisms in $w$ produce a difference at the origin $h$. If such $p$ exists, $w$ is nontrivial, as we can lift $p$ to $q\in A^{B_{(nr)^{t}}}$ without introducing forbidden patterns, and then $w$ acts nontrivially on the central $B_{nr}$-pattern of $q$, which by PEPB is a pattern appearing in $X$. On the other hand, if $w$ is nontrivial, then the algorithm succeeds in procing nontriviality since we can use $H=B_{(nr)^{t}}$ and any pattern $p$ where $w$ acts nontrivially.
Now suppose $G$ has EXPTIME word problem and $X$ has NP language. In this case, we can actually compute $B_{nr}$, and we can check whether a guessed pattern $p\in A^{B_{nr}}$ is in the language of $X$ by giving it as an input to the assumed NP algorithm. Since the pattern is of exponential size, the runtime is at most polynomial in $O(e^{n})$, thus we can do it in NEXPTIME. ∎
Appendix A Complexity theory
A.1 Basic complexity theoretic definitions
In this section, we give standard complexity theoretic definitions of Turing machines, P, NP, PSPACE, NEXPSPACE and NC1, as well as some terminology we use with them.
A complexity class is a set of languages $L\subset\Sigma^{*}$ for finite alphabets $\Sigma$.
A nondeterministic Turing machine $T$ is a tuple $(Q,\Sigma,\Delta,q_{\mathrm{init}},F,\square)$, where $Q$ is a finite set (of states), $\Sigma$ is an alphabet (the tape alphabet), $F\subset Q$ is the set of final states, $q_{\mathrm{init}}$ is the initial state, $\Delta\subset Q\times\Sigma\times Q\times\Sigma\times\{-1,0,1\}$, and $\square\in\Sigma$ is the blank symbol. An instantaneous description or ID is any element of $(\Sigma\cup(Q\times\Sigma))^{+}$ or $(\Sigma\cup(Q\times\Sigma))^{\omega}$ where exactly one symbol from $Q\times\Sigma$ appears. In the case of $\omega$, we require all but finitely many symbols are $\square$. The head position of an ID $w$ is the length of the unique word $u$ such that $w=u(q,s)v$ for some $q\in Q,s\in\Sigma$, and $q$ is the head state. It is sometimes useful to use another coding of IDs where we use words in $\Sigma^{*}Q\Sigma^{+}$, by using a word from $Q\Sigma$ in place of the $(Q,\Sigma)$-symbol.
An instantaneous description is initial if the head position is 00, and final if the head state is in $F$. Next we define computation steps (of type $(q,s,q^{\prime},s^{\prime},d)\in\Delta$). If $(q,s,q^{\prime},s^{\prime},0)\in\Delta$, then there is a computation step (of this type) from $u(q,s)v$ to $u(q^{\prime},s^{\prime})v$ for all (possibly infinite) words $u,v$ over $\Sigma$. If $(q,s,q^{\prime},s^{\prime},1)\in\Delta$, then there is a computation step from $u(q,s)av$ to $us^{\prime}(q^{\prime},a)v$ for all $u,v$, and any symbol $a\in\Sigma$. If $(q,s,q^{\prime},s^{\prime},-1)\in\Delta$, then there is a transition from $ua(q,s)v$ to $u(q^{\prime},a)s^{\prime}v$ for all $u,v$, and any symbol $a\in\Sigma$. A computation is a sequence of IDs where all adjacent IDs are computation steps. A computation is finite if it has finitely many IDs. The length of a finite computation is the number of IDs in it.
We say $T$ is deterministic if the implication $(q,s,q^{\prime},s^{\prime},d),(q,s,q^{\prime\prime},s^{\prime\prime},d^{\prime})\in\Delta\implies q^{\prime}=q^{\prime\prime},s^{\prime}=s^{\prime\prime},d=d^{\prime}$ holds for all $(q,s)\in Q\times S$. For $w=av\in\Sigma^{*}$ (with $a\in\Sigma$) write $q_{\mathrm{init}}w=(q_{\mathrm{init}},a)v$, and $q_{\mathrm{init}}\epsilon=q_{\mathrm{init}}\square$. We say $T$ accepts $w\in\Sigma^{*}$ in time $x\in\mathbb{N}$ and space $y\in\mathbb{N}$ if there is a computation whose initial ID is $q_{\mathrm{init}}w\square^{y-|w|}$ and length at most $x$. We may omit the time or space requirement separately.
Another less obvious coding we use is by Wang tiles. Specifically, there is a standard way to construct a finite set of Wang tiles $W\subset C^{\{\mathrm{north},\mathrm{south},\mathrm{west},\mathrm{east}\}}$ from a Turing machine, so that tilings of rectangles (satisfying the usual Wang tiling rule that the left (west) color of a tile must match the right (east) color of its left (west neighbor)) correspond to computations of the Turing machine. We omit the detailed construction (see e.g. [26]), but the idea is simply to use arrows to transmit the Turing machine head correctly (the colors ensure that the arrows match, and the head moves to the correct tile.
The complexity class P (resp. NP) contains the languages $L$ which are accepted by a deterministic (resp. nondeterministic) Turing machine in polynomial time. Specifically, $L\subset\Sigma^{*}$ is in NP if there exists a Turing machine $T$ and $d$ such that for all $w\in\Sigma^{*}$, $T$ accepts $w$ in polynomial time. The class PSPACE contains the languages accepted by a deterministic Turing machine with polynomial space (by a theorem of Savitch, one may equivalently use nondeterministic Turing machines). NEXPTIME contains the languages accepted in time $2^{O(n^{d})}$ for some $d$.
If $\mathcal{C}$ is a complexity class, co-$\mathcal{C}$ is the set of complements of languages in $\mathcal{C}$. We say a language $L\subset A^{*}$ is $\mathcal{C}$-hard if for each $L^{\prime}\in\mathcal{C}$ with $L^{\prime}\subset B^{*}$, there is a deterministic Turing machine which in polynomial time computes (in the sense that after it enters a final state, some specific track of its output contains the output word) from $w\in B^{*}$ a word $u\in A^{*}$ such that $u\in L$ if and only if $w\in L^{\prime}$. We say $L$ is $\mathcal{C}$-complete if it is in $\mathcal{C}$ and $\mathcal{C}$-hard.
[…]
Now, (nonuniform) NC1 (where NC stands for “Nick’s class”) is the class of languages $L\subset\{0,1\}^{*}$ such that for some $C$, for each $n$, there is a circuit of depth $C\log n$ that computes the characteristic function of $L\cap\{0,1\}^{n}$. (Often, polynomial size is also required, since it is relevant in many other circuit-based complexity classes, but in the present case, this is automatic from logarithmic depth.)
Lemma A.1.
In the definition of NC1, one may assume only $-ı\i$-labeled vertices of the circuit may have out-degree greater than $1$.
Proof.
The only requirement is logarithmic depth, so we may make as many copies of the nodes as we well, and there is no need to use the same node twice. Specifically, start from the output (nodes labeled by $o$), and traverse backward to the input, unfolding the acyclic graph to a tree (apart from incoming edges from the input nodes labeled by $-ı\i$). ∎
One may equivalently think of NC1 as the languages where for each length $n$, there is a Boolean formula describing the words of length $\{0,1\}^{n}$ (in terms of $n$ basic Boolean variables), and which has depth $O(\log n)$. Internal nodes with out-degree greater than $1$ correspond to variables being reused in this point of view, but as shown this is irrelevant for the strength of the model.
We recall a weak version of the Master Theorem from algorithm analysis: if a function $T(n)$ satisfies $T(n)\leq aT(n/b)+f(n)$, then writing $c=\log_{b}a$ and assuming that $f(n)$ is sandwiched between two polynomials, we have $T(n)=O(n^{c+1}+f(n))$. (We will not need a function $f(n)$.)
Proposition A.1.
The following is known about inclusions between the classes:
  • NC1 $\subset$ P $\subset$ NP $\subset$ PSPACE $\subset$ NEXPSPACE;
  • NC1 $\subsetneq$ PSPACE;
  • NP $\subsetneq$ NEXPSPACE.
A.2 Polynomial space and alternating polynomial time
Here, PSPACE is the class of problems that can be solved by a Turing machine in polynomial space (in the input length).
AP is the set of problems that can be solved by alternating Turing machines in polynomial time, meaning the Turing machine can take existential transitions (all states of a nondeterministic Turing machine can be thought of as being existential, and the word is accepted at a particular step if some transition leads to an accepting computation under the stated resource bounds) and universal transitions (in which case the word is accepted if all outgoing transitions lead to an accepting computation). We omit the precise definition, which is analogous to the definition of NP (or Lemma A.3 could be taken as a definition).
We will make strong use of the following classical result of Chandra, Kozen and Stockmeyer.
Lemma A.2 ([10]).
PSPACE = AP
Alternating polynomial time has a natural complete problem: Given a word $u$ of length $n$, does a fixed universal Turing machine $T$ accept $u$ in $n$ steps of computation?
Here, the universal machine $T$ is an alternating universal Turing machine that takes a description of any alternating Turing machine, and then simulates its computation, using the same alternation behavior as the simulated machine.
It is helpful in our applications to be a little more explicit about the universal quantification, as we will explicitly lay the tree of universal choices on the configuration, while nondeterministic steps will be nondeterministic (they will be coded in the configuration).
Lemma A.3.
There exist two nondeterministic Turing machines $T_{0},T_{1}$ sharing their initial state $q_{0}$, tape alphabet $S$ and state set $Q$ (but having possibly distinct nondeterministic rules), such that the following problem is PSPACE-hard: Given a word $u=av\in S^{n}$ with first letter $a\in S$ and with $n$ a power of $2$, does there exist a binary tree of Turing machine configurations $(c_{w})_{w\in\{0,1\}^{\leq n}}$ where $c_{\epsilon}=(a\times q_{0})v$, $(c_{w},c_{w0})$ is a step of $T_{0}$ and $(c_{w},c_{w1})$ is a step of $T_{1}$, and all $c_{w}$ with $|w|=n$ are accepting.
There are a few notes to make. Note that we are only using $|u|$ time in the computation, and the length of $u$ is a power of $2$. This is nothing to worry about, as in the definition of AP-hardness we ask for a polynomial time computable many-one reduction, and such a reduction is allowed to include a polynomial padding, meaning a polynomial time computation becomes linear. (If the computation does not end at the correct power of $2$ on some branch, the Turing machine can simply wait in the accepting state.)
The point of the two Turing machines is that we can use the steps of $T_{0}$ and $T_{1}$ as a universal quantification, as both branches must accept in all leaves. Of course, typically we do not want to universally quantify at every step, but we can simply have $T_{0}$ and $T_{1}$ behave the same, i.e. according to some alternating Turing machine $T$, except when $T$ makes a universal step, in which case $T_{0}$ and $T_{1}$ follow the two choices (or if there are more choices, this can be simulated by a bounded number of branchings).
A.3 (A version of) Barrington’s theorem
Barrington’s theorem [3] relates efficiently parallelizable computation (NC1) and computation with finite memory:
Theorem A.1 (Barrington’s theorem).
nonuniform NC1 = BWBP
We do not define BWBP, but we now define a “reversible” version BWBP’, for which a similar theorem holds.
First, we define a permutation group that essentially captures the type of computation performed by BWBP, namely the group of controlled permutations.
Definition A.1.
Let $A,B$ be finite sets and $n\in\mathbb{N}$. Define $G_{n,A,B}$ to be the permutation group acting on $B\times A^{n}$, generated by the following permutations:
  • $\sigma$ defined by $\sigma(b,x)=(b,\sigma^{\prime}(x))$, where $\sigma^{\prime}:A^{n}\to A^{n}$ is defined by $\sigma^{\prime}(au)=ua$ where $a\in A,u\in A^{n-1}$;
  • for each even permutation $\pi$ of $B$ and $a\in A$, the map $\pi|[a]$ defined by
    $\pi|[a](b,a^{\prime}u)=\begin{cases}(\pi(b),a^{\prime}u)&\mbox{ if}a^{\prime}=a,\\ (b,a^{\prime}u)&\mbox{ otherwise}\\ \end{cases}$
One should read $\pi|[a]$ as “perform permutation‘$\pi$ given that the input is in $[a]$”, or just “$\pi$ given $[a]$”, and the idea is that we perform permutation $\pi$ on the $B$-component if the $A^{n}$-component is in the cylinder $[a]=[a]_{0}=\{au\;|\;u\in A^{n-1}\}$.
We now set up the usual calculus of controlled permutations.
First we generalize the $\pi|[a]$ notation: for any set $X\subset A^{n}$ and $\pi\in\mathrm{Alt}(B)$ (even permutations of $B$), define
$\pi|U(b,x)=\begin{cases}(\pi(b),x)&\mbox{ if}x\in X,\\ (b,x)&\mbox{ otherwise}\\ \end{cases}.$
We refer to $A^{n}$-components of pairs $(b,u)\in B\times A^{n}$, or sets $X$ used in generators, as the control.
We will show that all these elements are in $G_{n,A,B}$ as soon as $|B|\geq 5$. For $u\in A^{*}$, define $[u]_{i}=\{w\in A^{n}\;|\;\forall j\in[0,|u|-1]:w_{i+j}=u_{j}\}$.
Lemma A.4.
For all $A,B,n$ we have $\pi|[a]_{i}\in G_{n,A,B}$.
Proof.
[…]
The generator $\sigma^{i}$ will only ever be used through this lemma. Note that $\pi|[a]_{i}$ has word norm $12i-$ 1. Thus, while technically the $A^{n}$-component is shuffled all the time in intermediate steps, one should think that the $A^{n}$-component always stays fixed and we are simply conditioning permutations $\pi$ on different positions. The reason we explicitly use the generator $\sigma$ is to have a uniform generating set, and so that $\pi|[a]_{i}$ has word norm $i$ (as it will in the applications).
Lemma A.5.
For all $X\subset A^{n}$ we have $\pi_{1}|X\circ\pi_{2}|X=(\pi_{1}\circ\pi_{2})|X$.
Proof.
The $A^{n}$-component is fixed by elements of the form $\pi|Y$, so the checks $u\in X$ preserve their truth value between applications. Thus if $u\in X$, a direct calculation shows
$(\pi_{1}|X\circ\pi_{2}|X)(b,u)=(\pi_{1}(\pi_{2}(b),u)=(\pi_{1}\circ\pi_{2})|X(b,u),$
and if $u\notin X$, then all of $\pi_{1}|X,\pi_{2}|X,(\pi_{1}\circ\pi_{2})|X$ act trivially. ∎
Lemma A.6.
For all $X\subset A^{n}$ we have $\pi^{-1}|X=(\pi|X)^{-1}$
Proof.
Since $X$ is finite, it suffices to check that $\pi^{-1}|X$ is a one-sided inverse of $\pi|X$. Indeed we have $\pi^{-1}|X\circ\pi|X=(\pi^{-1}\circ\pi)|X=\mathrm{id}|X=\mathrm{id}$ by the previous lemma. ∎
We will often use the following lemma about the alternating group (it expresses that the commutator width of the group is one).
Lemma A.7.
Let $|B|\geq 5$ be a finite set. Then for all $\pi\in\mathrm{Alt}(B)$ we have $\pi=[\pi_{1},\pi_{2}]$ for some $\pi_{1},\pi_{2}\in\mathrm{Alt}(B)$.
Remark A.1.
The statement of this lemma is nice and clean, but a few remarks on its necessity may be in order, since we are not aware of a proof that avoids case distinctions. The lemma was stated by Ore in [24], and proved by Ito in [20]. Ito’s paper seems difficult to find, but a proof (of a stronger result about even permutations) is found also in [5]. It is known that the commutator width is one in all simple groups. We need the lemma only for a fixed set of cardinalities $|B|$, so one could simply check it by computer. It even suffices that the commutator width if bounded (over all $|B|$), although this changes the degrees of polynomial bounds. One can even phrase the argument directly in terms of $3$-rotations to avoid using the lemma entirely by using simply $[(a\;b\;c),(c\;d\;e)]=(c,b,e)$; this latter is the approach taken by Barrington. We also sometimes use only $3$-rotations, namely in some cases we need the support to be non-full.
Lemma A.8.
For all $A,B,n$ and $X,Y\subset A^{n}$ we have the equality $[\pi_{1}|X,\pi_{2}|Y]=[\pi_{1},\pi_{2}]|X\cap Y$.
Proof.
If $u\notin X$ then
$[\pi_{1}|X,\pi_{2}|Y](b,u)$ $=(\pi_{1}^{-1}|X\circ\pi_{2}^{-1}|Y\circ\pi_{1}|X\circ\pi_{2}|Y)(b,u)$
$=(\pi_{2}^{-1}|Y\circ\pi_{2}|Y)(b,u)$
$=(b,u)=[\pi,\pi^{\prime}]|X\cap Y(b,u)$
where the second equality follows because the $A^{n}$-component $u$ is never modified ($\sigma$ is only applied in intermediate steps). A similar calculation works if $u\notin Y$. If $u\in X\cap Y$ then clearly
$[\pi_{1}|X,\pi_{2}|Y](b,u)=([\pi_{1},\pi_{2}](b),u)=([\pi_{1},\pi_{2}]|X\cap Y)(b,u)$
since all of $u\in X,u\in Y,u\in X\cap Y$ hold at every step. ∎
The following is an immediate induction.
Lemma A.9.
$[\pi_{1}|X_{1},\pi_{2}|X_{2},\ldots,\pi_{k}|X_{k}]=[\pi_{1},\pi_{2},\ldots,\pi_{k}]|\bigcap_{i}X_{i}$
We note that for complexity-theoretic purposes, this is an inefficient definition, and one should balance the brackets instead. However, we only do this for commutators of bounded length, and this simple definition sufficies.
Lemma A.10.
For all $A,B,n$ and $X,Y\subset A^{n}$ we have
$\pi|\bar{X}=\pi\circ\pi^{-1}|X$
and
$\pi|X\cup Y=\pi|X\circ\pi|Y\circ\pi^{-1}|X\cap Y.$
Proof.
We give semantic proofs. In the first formula, on the right side $\pi$ gets applied anyway, but this cancels when also $\pi^{-1}|X$ applies. Thus, $\pi$ is ultimately applied if and only if the control is not in $X$.
In the second formula, $\pi|X\circ\pi|Y$ applies $\pi$ if the control is in $X$ or $Y$. The only problem is it applies it twice if it is in $X\cap Y$, and this is fixed by also applying $\pi^{-1}|X\cap Y$. ∎
Lemma A.11.
For all $A,n$ and $X\subset A^{n}$, if $|B|\geq 5$ then $\pi|X\in G_{n,A,B}$.
Proof.
Let $X\subset A^{n}$ and enumerate $X=\{u^{1},u^{2},\ldots,u^{k}\}$ so that $X=\bigcup_{i=1}^{k}[u^{i}]_{0}$. Then $\pi|X=\prod_{i}\pi|[u^{i}]_{0}$. On the other hand for each $u=u_{0}u_{1}\cdots u_{n-1}\in A^{n}$, letting $\pi=[\pi_{0},\pi_{1},\ldots,\pi_{n-1}]$ we have
$\pi|[u]_{0}$ $=[\pi_{0},\pi_{1},\ldots,\pi_{n-1}]|\bigcap_{i=0}^{n-1}[u_{i}]_{i}$
$=[\pi_{0}|[u_{0}]_{0},\pi_{1}|[u_{1}]_{1},\ldots,\pi_{n-1}|[u_{n-1}]_{n-1}]\in G_{n,A,B}$
by Lemma A.9. ∎
The lengths of some controlled permutations are necessarily very long, since the group is very large. BWBP’ is defined through controlled permutations with small word norm.
Definition A.2.
A language $L\subset A^{*}$ is in BWBP’B if for some $d$, for all $n$ and $\pi\in\mathrm{Alt}(B)$, the norm of $\pi|L\cap A^{n}$ in $G_{n,A,B}$ is at most $n^{d}+d$. Write
$\mathrm{BWBP^{\prime}}=\bigcup_{B\;\mathrm{finite}}\mathrm{BWBP^{\prime}}_{B}.$
(Writing $n^{d}+d$ here is equivalent to writing $O(n^{d})$.)
We now prove the relevant part of Barrington’s theorem.
Theorem A.2 (variant of Barrington’s theorem).
nonuniform NC1 $\subset$ BWBP’B for all $|B|\geq 5$
Proof.
Let $A=\{0,1\}$. Let $B$ be any alphabet with cardinality at least $5$. Let $s=(s_{1},s_{2}):\mathrm{Alt}(B)\to\mathrm{Alt}(B)\times\mathrm{Alt}(B)$ be any function such that $\pi=[s_{1}(\pi),s_{2}(\pi)$ for all $\pi$, which exists by Lemma A.7. Identify Boolean formulas $\phi$ with the subsets of $A^{n}$ they describe. Then by Lemma A.8 and Lemma A.10 we have
$\pi|\phi_{1}\wedge\phi_{2}=[s_{1}(\pi)|\phi_{1},s_{2}(\pi)|\phi_{2}]$
$\pi|\phi_{1}\vee\phi_{2}=\pi|\phi_{1}\circ\pi|\phi_{2}\circ[s_{1}(\pi)^{-1}|\phi_{1},s_{2}(\pi)^{-1}|\phi_{2}]$
$\pi|\neq\phi=\pi\circ\pi^{-1}|\phi$
Let $T_{d}$ be the maximal word norm of $\pi|\phi$ for $\phi$ of depth $k=O(\log n)$, when using generators of the form $\pi|[a]_{i}$ for $a\in A$. The formulas show $T_{i}\leq 6T_{i-1}$ for all $i$, so $T_{k}=O(6^{k})=O(6^{\log n})=O(n^{d})$ for some $d$. Finally, each $\pi|[a]_{i}$ has word norm $O(n)$. ∎
We now give the main examples of languages are in BWBP’.
For the next lemma, let $S$ be the tape alphabet of a Turing machine $T$, and $Q$ its state set. Let $C_{S,Q}\subset((S\times Q)\cup S)^{*}$ be the language of words where a single $S\times Q$-symbol appears. In other words, the language of Turing machine configurations.
Lemma A.12.
The language $L=\{u\#v\;|\;(u,v)\mbox{ is a computation step }\}\subset((S\times Q)\cup S\cup\{\#\})^{*}$ is in BWBP’, for any Turing machine $T$ with tape alphabet $S$ and state set $Q$.
Proof.
Let $A=((S\times Q)\cup S\cup\{\#\})^{*}$. We show that a logarithmic-depth formula exists for inputs $w=uav$ of length $12n+$ 1 with $u\in A^{n},a\in A,v\in A^{n}$, i.e. that the language is in NC1. Note that Boolean formulas can only talk about bits, but we can take the basic Boolean variables to be statements of the form “$w_{i}=a$” for $i\in[0,2n]$. For $I\subset\{0,\ldots,2n+1\}$, we use names of the form $\texttt{FORMULA}_{I}(x_{1},\ldots,x_{k})$ for Boolean formulas. Formally, this is just a name, but the label FORMULA and the parameters $(x_{1},\ldots,x_{k})$ explain what the formula is describing, and the subscript $I$ means that the formula ultimately only refers to coordinates $I$ of the input $w$. Omitting the subscript means $I=[0,2n]$.
The point of $I$ is that logarithmic depth is obvious as soon as we always roughly cut the size of $I$ in half. We will describe the bottom of the recursion, and only give the formulas that apply for large $I$, since for small $I$ (say $|I|\leq 4$, there exists a formula of constant size by the simple fact that Boolean formulas can describe any set.
First, we define formulas checking $a=\#$ (i.e. the $n$th coordinate of $w$ contains ##\##) and #⊏̸$\#\not\sqsubset uv# ⊏$. For this, define $\texttt{SUBALPHABET}_{I}(B)$ to be the check that $w_{i}\in B$ for all $i\in I$. The formula $\texttt{SUBALPHABET}_{I}(B)=\texttt{SUBALPHABET}_{J}(B)\wedge\texttt{SUBALPHABET}_{K}(B)$ shows that this formula always has logarithmic depth, by using cuts $I=J\sqcup K$ with $|J|\approx|K|\approx|I|/2$. Now the desired formula is
$\texttt{HASHCHECK}()=\texttt{ALPHABET}_{\{n\}}(\{\#\})\wedge\texttt{ALPHABET}_{[0,n-1]\cup[n+1,2n]}(A\setminus\{\#\}).$
Next, we check the main part, i.e. that a single Turing machine step is being described. We write $\texttt{STEP}()$ for this formula, where $I$. It suffices to show for each $(q,s,q^{\prime},s^{\prime},d)\in\Delta$ we have a logarithmic depth formula for $\texttt{STEP}(q,s,q^{\prime},s^{\prime},d)$, which accepts precisely the steps of type $(q,s,q^{\prime},s^{\prime},d)$. Namely then
$\texttt{STEP}()=\bigvee_{(q,s,q^{\prime},s^{\prime},d)\in\Delta}\texttt{STEP}(q,s,q^{\prime},s^{\prime},d).$
We only show how to define $\texttt{STEP}(q,s,q^{\prime},s^{\prime},1)$, as $\texttt{STEP}(q,s,q^{\prime},s^{\prime},0)$ is strictly easier and $\texttt{STEP}(q,s,q^{\prime},s^{\prime},-1)$ is completely symmetric.
We need some auxiliary firmulas. As a first auxiliary formula, we define $\texttt{BASICSTEPAT}_{\{i,i+1,i+n+1,i+n+2\}}(i,q,s,q^{\prime},s^{\prime},1)$ as the check that
$w_{i}w_{i+1}w_{i+n+1}w_{i+n+2}=((q,s),a,s^{\prime},(q^{\prime},a^{\prime}))$
for some $a,a^{\prime}\in S$. Since this formula talks about only $4$ positions of the input, it has a formula of constant size.
As another auxiliary formula, for $J,K\subset[0,2n+1]$ with $|J|=|K|$, define $\texttt{EQ}_{J\sqcup K}(J,K)$ as the check that $w_{j}=w_{k}$ when $j$ and $k$ are the $i$th smallest element of $J,K$ respectively (i.e. the subsequences along $J$ and $K$ are equal). We have
$\texttt{EQ}_{J\cup K}(J,K)=\texttt{EQ}_{J^{\prime}\cup K^{\prime}}(J^{\prime},K^{\prime})\wedge\texttt{EQ}_{J^{\prime\prime}\cup K^{\prime\prime}}(J^{\prime\prime},K^{\prime\prime}),$
where $J=J^{\prime}\sqcup J^{\prime\prime}$ and $K=K^{\prime}\sqcup K^{\prime\prime}$ are roughly equal cuts at the middle with $|J^{\prime}|=|K^{\prime}|$. Then clearly $\texttt{EQ}_{J\cup K}(J,K)$ always has a logarithmic depth formula.
We now observe that
$\texttt{STEP}(q,s,q^{\prime},s^{\prime},1)=\bigvee_{i\in[0,n-2]}($ $\texttt{BASICSTEP}_{{i,i+1,i+n+1,i+n+2}}(i,q,s,q^{\prime},s^{\prime},1)$
$\wedge\texttt{EQ}_{J_{i}\cup K_{i}}(J_{i},K_{i}))$
where $J_{i}=[0,n-1]\setminus\{i,i+1\},K_{i}=[n+1,2n]\setminus\{i+n+1,i+n+2\}$. Finally, a logarithmic depth formula for a polynomial-sized disjunction $\bigvee_{i\in[0,n-2]}\phi_{i}$ is obtained by associating it in a balanced way.
We conclude that there is a formula
$\texttt{L}()=\texttt{STEP}()\wedge\texttt{HASHCHECK}()$
of logarithmic depth in $n$, which describes $L$.
By Theorem A.2, $\pi|L$ has polynomial norm in $G_{n,A,B}$ for $|B|\geq 5$. ∎
Having written one proof in detail, we will from now on not explicitly write the formulas, and only explain why a divide-and-conquer approach is possible. In each case, we can split the problem into a bounded number of subproblems, where the number of bits being looked at becomes roughly cut in half.
Lemma A.13.
For any sequence of words $u_{i}\in\Sigma^{i}$, the language $\{u_{i}\;|\;i\in\mathbb{N}\}$ is in BWBP’.
Proof.
To compare $w\in\Sigma^{i}$ with $u_{i}$, split $w$ in half, and recursively compare the left and right halves. ∎
For the next lemma for $n\in[0,m]$, denote by $\mathrm{unary}_{m}(n)=0^{n}1^{m-n}$ the length-$m$ unary representation of $n$.
Lemma A.14.
The language $\{\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)\;|\;n\in[0,m)\}\subset\{0,1,\#\}^{*}$ is in BWBP’. Similarly, the relation $\{u2^{n}\#ua2^{n}\;|\;u\in\{0,1\}^{*},n\in\mathbb{N}\}$ is in BWBP’ for each $a\in\{0,1\}$.
Proof.
For simplicity assume $m$ is a power of $2$. For the divide-and-conquer, observe that checking the input to be in the form $\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)$ can be split in two cases after splitting the input in half:
  • the left halves are all zero, and the right halves are in the increment relation;
  • the right halves are all one, and the left halves are in the increment relation.
The other relation is shown to be in BWBP’ by first checking that the projection $0,1\mapsto 0;2\mapsto 1$ give a word in the first language, then checking the binary symbols are in the same position, or in case only the second word carries a bit, this bit is $a$. (This latter is a matter of checking that all pairs of corresponding positions in the two parts belong to a particular set.) ∎
The increment relation for binary numbers is also in BWBP’. Say numbers are written least significant bit first and 00-indexed, i.e. define $\mathrm{val}(u)=\sum_{i=0}^{|u|-1}2^{i}u_{i}$ for a word $u\in\{0,1\}^{*}$. Let $\mathrm{bin}_{m}:[0,2^{m})\to\{0,1\}^{m}$ be the inverse of this function, i.e. it gives the binary representation of a natural number.
Lemma A.15.
The language $\{\mathrm{bin}_{m}(n)\#\mathrm{bin}_{m}(n+1)\;|\;n\in[0,2^{m})\}\subset\{0,1,\#\}^{*}$ is in BWBP’.
Proof.
Again for simplicity assume $m$ is a power of $2$. For the divide-and-conquer, observe that $\mathrm{bin}_{m}(n)\#\mathrm{bin}_{m}(n+1)$ can be split in finitely many cases:
  • the left halves are equal, and the right halves are in the increment relation
  • the right halves are respectively $1^{m/2}$ and $0^{m/2}$, and the left halves are in the increment relation. ∎
More generally, checking for a fixed constant sum can be performed in polynomial word norm.
Lemma A.16.
Fix $d$, $|B|\geq 5$ and any $\pi\in\mathrm{Alt}(B)$. Then for any $\ell$, the permutation $\pi|\{\mathrm{bin}_{m}(n_{1})\#\mathrm{bin}_{m}(n_{2})\#\cdots\#\mathrm{bin}_{m}(n_{d})\;|\;\sum_{i}n_{i}=\ell\}$ has word norm polynomial in $m$ in $\mathcal{G}_{dm+d-1,A,B}$.
Proof.
It suffices to construct a circuit whose output is the sum of two numbers, as one can chain $d-1$ of these together and finally compare the sum to a fixed number using Lemma A.13.
This is [1, Example 6.19]. The bits of the sum are determined by the added bits at that position, as well as the carry signal coming from the left (from the left, since we use LSB for binary numbers). It is easy to show that the carry signal at a particular position can in turn be “predicted” by a constant-depth circuit, when we allow conjunctions and disjunctions (circuit nodes labeled $\wedge$ and $\vee$) with arbitrarily many inputs. Associating these in a balanced way gives a logarithmic depth circuit where all nodes have in-degree $2$. ∎
Remark A.2.
In modern CPU design this corresponds roughly to the carry-lookahead adder construction (speeding up the school algorithm of linear carry propagation, which in turn is known as a ripple-carry adder in circuit design); Babbage invented a similar anticipating carriage for his analytical engine.